### v1

# TAR reconfiguration for vertex set parameters

Abstract
This paper surveys results about token addition and removal (TAR) reconfiguration for several well-known vertex set parameters including domination, power domination, standard zero forcing, and PSD zero forcing. We also expand the range of parameters to which universal $X$-set TAR graph results apply, for $X$-sets and their TAR graphs introduced in [B. Bjorkman, C. Bozeman, D. Ferrero, M. Flagg, C. Grood, L. Hogben, B. Jacob, C. Reinhart, Power domination reconfiguration, https://arxiv.org/abs/2201.01798] and [N.H. Bong, J. Carlson, B. Curtis, R. Haas, L. Hogben, Isomorphisms and properties of TAR reconfiguration graphs for zero forcing and other $X$-set parameters, Graphs Combin. 39 (2023), Paper No. 86]. Here we examine which of the $X$-set axioms are needed for which results. With this new framework, the main results apply to skew zero forcing and vertex covering, and results about TAR reconfiguration graphs of these parameters are presented. While $X$-sets are defined for parameters that take the minimum cardinality over the $X$-sets of a graph, and $X$-set results are restricted to such minimizing parameters, our expansion of the universal perspective allows these results to be applied to parameters that take the maximum value among relevant sets, called $Y$-sets. Maximizing parameters to which the main results apply include independence number, (upper) irredundance number, and (upper) zero forcing irredundance number, and failed zero forcing number; TAR reconfiguration results are presented for these parameters. We also show that the equivalence of connectedness in certain token jumping reconfiguration graphs and certain TAR reconfiguration graphs for independent sets established in [M. Kamiński, P. Medvedev, M. Milanič. Complexity of independent set reconfiguration problems. J. Theoretical Computer Science 439 (2012), 9–15.] extends to $X$-set and $Y$-set parameters.
Keywords reconfiguration; token addition and removal; TAR; vertex set; $X$-set; $Y$-set; domination; power domination; zero forcing; PSD forcing; skew forcing; vertex cover; independence number; irredundance; failed zero forcing
AMS subject classification 05C69, 05C45, 05C50, 05C57, 05C60, 05C70, 68R10
1 Introduction
The study of reconfiguration examines relationships among solutions to a problem. The reconfiguration graph for the problem has as its vertices solutions to a problem and edges are determined by a reconfiguration rule that describes relationships between the solutions. The reconfiguration rule can be viewed as describing a single step in the process of transforming one solution to another in which each intermediate state is also a solution. The ability to transform one solution to another solution is equivalent to having a path between the two solutions in the reconfiguration graph, i.e., the two solutions are in the same connected component of the reconfiguration graph. In [34], Nishimura surveys recent work on structural and algorithmic (complexity) questions across a broad variety of parameters. She describes three types of reconfiguration rules, including token addition and removal.
For many reconfiguration problems arising from graphs (including all those we study here), a solution can be represented as a subset of the vertices of a graph $G$; we call such a problem a vertex set problem. The token addition and removal (TAR) reconfiguration graph for a vertex set problem has an edge between two sets if and only if one can be obtained from the other by the addition or removal of a single vertex. Both the TAR graph (which includes all feasible subsets of vertices of $G$) and the $k$-TAR graph, which allows only feasible subsets of at most (or at least) $k$ vertices, are studied.
A set $S\subseteq V(G)$ is a dominating set of a graph $G$ if every vertex of $G$ is in $S$ or a neighbor of a vertex in $S$. The domination number of $G$, denoted by $\gamma(G)$, is the minimum cardinality of a dominating set of $G$. The domination number has been widely studied; see for example, the book by Haynes, Hedetniemi, and Slater [23] and more recent books by Haynes, Hedetniemi, and Henning, such as [22]. Haas and Seyffarth initiated the study of reconfiguration graphs for domination in [19]. Many connectedness results for $X$-set parameters in [6] use the same ideas as the analogous results for domination.
A universal definition of an $X$-set for a parameter determined by subsets of vertices was given in [6] and [7]; parameters to which this definition applies are here called original $X$-set parameters and include the domination number, the power domination number, the standard zero forcing number, and the PSD zero forcing number, but not to the skew forcing number (see Section 1.2 for definitions of these parameters). Bjorkman et al.  were the first to use the universal approach with $X$-sets. In addition to establishing connectedness results, they showed that for an on original $X$-set parameter and a graph $G$ with no isolated vertices, the order of $G$ can be determined from its $X$-TAR graph. Power domination TAR graphs were also studied in detail in that paper (as were token jumping reconfiguration graphs, also known as token exchange). In [7] numerous additional results about $X$-TAR graphs were established, including that for an on original $X$-set parameter and graphs $G$ and $G^{\prime}$ with no isolated vertices, an isomorphism of the $X$-TAR graphs can be used to find a relabeling of the vertices of $G^{\prime}$ so that $G$ and $G^{\prime}$ have the same $X$-sets; standard zero forcing TAR graphs were also studied in that paper.
Skew zero forcing is an obvious omission from the list of parameters to which the results in [6] and [7] apply, because skew zero forcing allows the empty set to be a skew forcing set, and this was prohibited by one of the rules for $X$-sets in [6, 7] (see Definition 2.1). In Section 2 we determine which of the axioms in the original definition of $X$-set are needed for each result of [6] and [7]. We present both a very general definition of a super $X$-set parameter (Definition 2.4), which uses only one of the original $X$-set parameter axioms, and a stronger robust $X$-set parameter definition (Definition 2.5), which uses versions of three of the five axioms for original $X$-set parameters. Many connectedness results stated in [6] (and derived from earlier work such as [19]) and a result from [30] about token jumping graphs and TAR graphs are established for every super $X$-set parameter, but the isomorphism results need not be true for such a parameter. Almost all results of [6, 7] are true for every robust $X$-set parameter, allowing the extensions of these results to additional parameters including skew forcing number and vertex cover number.
In Section 3, we survey known results about TAR graphs for various parameters related to domination and zero forcing, primarily about connectedness of TAR subgraphs, isomorphism properties, uniqueness, and existence of Hamilton cycles or paths. Parameters that have been studied include domination number, power domination number, and standard zero forcing number; we also present a small additional amount of material for these parameters. TAR graphs for PSD forcing and skew forcing are examined here in more detail in Sections 4 and 5, where examples illustrating connectedness, uniqueness, and Hamiltonicity are presented, in addition to the main order and isomorphism results and connectedness bounds that follow from the revised $X$-set parameter definitions. Reconfiguration of vertex covers was studied in [28, 29] and other works. Complexity is the focus of much of this prior work on vertex cover reconfiguration, but some structural results have been obtained and we provide a partial survey of these results in Section 6. We do not survey complexity results for any of the parameters discussed.
We initially take a very general approach in Section 2 by defining a vertex set parameter, which allows the main results to be extended to many vertex set parameters that take the maximum size of a set rather than the minimum; such parameters are here called $Y$-set parameters. General results for TAR graphs of $Y$-set parameters are discussed in Section 8, together with a more detailed discussion on TAR graphs for independence, irredunundance, zero forcing irredundance, and failed zero forcing. Two complementation techniques are introduced to translate $X$-set parameter results to $Y$-set parameter results.
In this arXiv version we have included the proofs of some results from other papers when the result is not stated as generally as it is here but essentially the same proof works; these are colored blue. In the next section we list notation we will use throughout. Section 1.2 provides definitions of the parameters that are discussed.
1.1 Notation and terminology
Every graph discussed is simple, undirected, finite, and has a nonempty vertex set. A graph $G=(V(G),E(G))$ consists of a set of vertices $V(G)$ and a set of edges $E(G)$ which are two element subsets of $V(G)$. The edge $\{v,w\}\in E(G)$ is denoted $vw$. The order of the graph is $|V(G)|$. A graph is odd or even according as its order is odd or even. The subgraph of $G$ induced by a set $W$ of vertices is denoted by $G[W]$. For $S\subseteq V(G)$, define $G-S=G[V(G)\setminus S]$. When $S$ is a single vertex $v\in V(G)$, the notation $G-\{v\}$ is simplified to $G-v$.
If $uw\in E(G)$, then $v$ and $w$ are said to be adjacent or neighbors. The neighborhood of $v$ is the set of neighbors of $v$ and is denoted by $N_{G}(v)$ and the closed neighborhood of $v$ is $N_{G}[v]=N_{G}(v)\cup\{v\}$. The closed neighborhood of a set $S\subset V(G)$ is $N_{G}[S]=\cup_{x\in S}N_{G}[x]$. The subscript $G$ may be omitted when it is clear from context. The degree of the vertex $v$ is $\deg_{G}(v)=|N(v)|$. For a graph $G$, $\Delta(G)=\max\{\deg_{G}(v):v\in V(G)\}$ and $\delta(G)=\min\{\deg_{G}(v):v\in V(G)\}$.
Given distinct vertices $v_{0}$ and $v_{\ell}$ in $V(G)$, a path of length $\ell$ from $v_{0}$ to $v_{\ell}$ is a sequence of distinct vertices $(v_{0},\ldots,v_{\ell})$ such that $v_{i}$ is a neighbor of $v_{i+1}$ for every integer $i$, $10\leq i\leq\ell-$ 1. A cycle on $k$ vertices, denoted $(c_{1},\dots,c_{k})$, is a sequence of distinct vertices $c_{1},\dots,c_{k}$ with an edge between $c_{i}$ and $c_{i+1}$ for all $11\leq i\leq k-$ 1 and the edge $c_{1}c_{k}$. A Hamilton cycle (respectively, Hamilton path) of a graph $G$ is a cycle (respectively, path) that includes all vertices of $G$.
A graph $G$ is connected if there exists a path between any two distinct vertices of $G$. If $u$ and $v$ are distinct vertices in a connected graph $G$, the distance between $u$ and $v$, denoted by $\operatorname{dist}_{G}(u,v)$, is defined to be the minimum length over all paths between $u$ and $v$. The diameter of a connected graph $G$ is the maximum value of $\operatorname{dist}_{G}(u,v)$ over all pairs of distinct vertices $u$ and $v$ of $G$ and is denoted by $\operatorname{diam}(G)$. The connected components of a graph $G$ are the maximal connected induced subgraphs of $G$. Two graphs $G$ and $H$ are disjoint if $V(G)\cap V(H)=\emptyset$ and the disjoint union of $G$ and $H$ is denoted by $G\sqcup H$. The symmetric difference of sets $A$ and $B$ is denoted by $A\ominus B=(A\cup B)\setminus(A\cap B)$.
Suppose $G_{1}$ and $G_{2}$ are disjoint graphs. The Cartesian product of $G_{1}$ and $G_{2}$, denoted by $G_{1}\square G_{2}$, is the graph with $V(G\square G_{2})=V(G_{1})\times V(G_{2})$ such that $(v_{1},v_{2})$ and $(u_{1},u_{2})$ are adjacent if and only if $v_{1}=u_{1}$ and $v_{2}u_{2}\in E(G_{2})$, or $v_{2}=u_{2}$ and $v_{1}u_{1}\in E(G_{1})$. The join of $G_{1}$ and $G_{2}$ is the graph $G_{1}\vee G_{2}$ with vertex set $V(G_{1}\vee G_{2})=V(G_{1})\cup V(G_{2})$ and edge set $E(G_{1})\cup E(G_{2})\cup\{v_{1}v_{2}:v_{1}\in V(G_{1}),v_{2}\in V(G_{2})\}$.
Given an integer $n\geq 1$, the path, cycle and complete graph on $n$ vertices are denoted by $P_{n}$, $C_{n}$ and $K_{n}$, respectively. In examples, the vertices of $P_{n}$ and $C_{n}$ will be labeled with the integers $1,\dots,n$, and path (cycle) order refers to the convention that the sequence of vertices in the path (cycle) is $(1,2,\dots,n)$. The empty graph $\overline{K_{n}}$ is the graph with $n$ vertices and no edges. We denote the the complete bipartite graph by $K_{p,q}$ and assume $1\leq p\leq q$ and the partite sets are $A=\{a_{1},\dots,a_{p}\}$ and $B=\{b_{1},\dots,b_{q}\}$.
1.2 Parameters
In this section we define most of the parameters we discuss. Standard zero forcing was introduced in multiple applications, including control of quantum systems and as an upper bound to the maximum nullity of a real symmetric matrix $A$ whose off-diagonal entries $a_{ij}$ are nonzero if and only if the graph has edge $ij$. Other types of zero forcing were defined to serve as similar bounds for maximum nullity among positive semidefinite or skew-symmetric matrices (the distinguishing factor for skew zero forcing is that the diagonal entries must all be zero).
Starting with an initial set of blue vertices $S$, a zero forcing process colors vertices blue by repeated applications of a color change rule; the color change rule determines the type of zero forcing. Here we discuss three types of zero forcing: standard, introduced in [3]; positive semidefinite (PSD), introduced in [5]; and skew, introduced in [27]. In each case, $B$ denotes the set of (currently) blue vertices and the set of (currently) white vertices $W=V(G)\setminus B$.
  • Standard color change rule: Any blue vertex $u\in B$ can change the color of a white vertex $w$ to blue if $w$ is the only white neighbor of $u$, i.e., $N(u)\cap W=\{w\}$.
  • PSD color change rule: Let $W_{1},\dots,W_{k}$ be the sets of vertices of the $k\geq 1$ components of $G[W]$. If $u\in B$, $w\in W_{i}$, and $w$ is the only white neighbor of $u$ in $G[W_{i}\cup B]$, then $u$ can change the color $w$ to blue.
  • Skew color change rule: Any vertex $u\in V(G)$ can change the color of a white vertex $w$ to blue if $w$ is the only white neighbor of $u$, i.e., $N(u)\cap W=\{w\}$.
Given an initial set $S$ of blue vertices of $G$, the standard, PSD, or skew final coloring is the subset of vertices that are blue after applying the relevant color change rule until no more changes are possible (the final coloring is unique for each of these color change rules). A standard, PSD, or skew zero forcing set for $G$ is a subset $S$ of vertices $S$ such that the final coloring is $V(G)$ when starting with exactly the vertices in $S$ blue. The standard, PSD, or skew zero forcing number of a graph $G$ is the minimum cardinality of a standard, PSD, or skew zero forcing set; this parameter is denoted by $\operatorname{Z}(G)$, $\operatorname{Z}_{+}(G)$, or $\operatorname{Z}_{-}(G)$, respectively. All these parameters (and power domination) are also discussed in [25, Chapter 9]. TAR reconfiguration graphs of standard zero forcing were studied in [7], where it was noted that the main results also apply to positive semidefinite TAR reconfiguration graphs; these parameters are discussed here in Sections 3.3 and 4, respectively. TAR reconfiguration graphs of skew zero forcing has not been studied previously and is discussed here in Section 5.
A set $S\subseteq V(G)$ is a power dominating set of $G$ if $N[S]$ is a standard zero forcing set of $G$. Power domination on graphs was introduced to model the placement of Phase Monitoring Unities (PMUs), which are used to monitor electric networks to avoid catastrophic failures (see, for example, [21] and [8]). TAR reconfiguration graphs of power domination were studied in [6] and are discussed Section 3.2.
The study of failed (standard) zero zero forcing was introduced by Fetcie, Jacob, and Saavedra in [16]. Other authors have studied failed skew forcing in [4] and failed power domination in [18]. A set $T\subseteq V(G)$ is a failed standard zero forcing set if the final coloring is not $V(G)$ when starting with exactly the vertices in $T$ blue. That is, $T$ is a failed standard zero forcing set if and only if $T$ is not a standard zero forcing set for $G$. TAR reconfiguration graphs of failed standard zero forcing are discussed Section 8.6.
[…]
A set $T$ of vertices in a graph $G$ is independent (or is an independent set) of $G$ if no two distinct vertices in $T$ are adjacent. The independence number of $G$, denoted by $\alpha(G)$, is the maximum cardinality of an independent set of vertices of $G$. A set $S$ of vertices in a graph $G$ is a vertex cover of $G$ if every edge of $G$ has at least one of its endpoints in $S$. The vertex cover number of $G$, denoted by $\tau(G)$, is the minimum cardinality of a vertex cover of vertices of $G$. The independence number and vertex cover number are widely studied parameters. It is well-known (and easy to see) that a set $T\subseteq V(G)$ is independent if and only if $V(G)\setminus T$ is a vertex cover of $G$. TAR reconfiguration graphs of vertex cover number and independence number are discussed Sections 6 and 8.3.
2 $X$-set parameters (supersets and minimal sets): Axioms required for specific results
Numerous results for $X$-set parameters are established in [6] and [7]. Many of these results still hold when some of the axioms in the definition of an $X$-set are removed. In this section we examine which of these axioms are necessary for each of the aforementioned results. This is accomplished by introducing several universal variants of the original definition of $X$-sets, including the definition of robust $X$-set parameters for which the main results of [6, 7] still hold. A benefit of this definition is that these results now apply to skew forcing, which had been missing from the family of parameters related to zero forcing. By allowing consideration of properties that do not require every isolated vertex to be in an $X$-set, these results apply to vertex covers (without any assumption about isolated vertices). Just as [6] extended some results on domination to original $X$-set parameters, later in this section we also extend additional results from domination to super $X$-set parameters (sometimes requiring additional axioms), including results on cut-vertices and Hamilton paths.
The next definition states the axioms that were assumed (including implicit assumptions) for $X$-set parameters as defined in Definitions 2.1 in [6] and 1.2 in [7].
Definition 2.1.
An original $X$-set parameter is a graph parameter $X(G)$ defined to be the minimum cardinality of an $X$-set of $G$, where the $X$-sets of $G$ are defined by a given property and satisfy the following axioms.
  1. (i)
    (Superset) If $S$ is an $X$-set and $S\subseteq S^{\prime}$, then $S^{\prime}$ is an $X$-set.
  2. (ii)
    ($(n-1)$-set) If $G$ has no isolated vertices, then every set of $|V(G)|-1$ vertices is an $X$-set.
  3. (iii)
    (Component consistency) Suppose $G=G_{1}\,\sqcup\,\dots\,\sqcup\,G_{k}$ where $G_{i},i=1,\dots,k$ are the connected components of $G$. Then $S$ is an $X$-set of $G$ if and only if $S\cap V(G_{i})$ is an $X$-set of $G_{i}$ for $i=1,\dots,k$.
  4. (iv)
    (Isolated vertex) $X(K_{1})=1$.
  5. (v)
    (Empty set) The empty set is not an $X$-set.
Note that the Isolated vertex axiom (iv) or similar was clearly intended in [6, 7] but not stated; without this or another assumption about $K_{1}$ having an $X$-set, a graph with no edges might have no $X$-sets. The Connected component axiom (iii) was stated only for one direction: If $S$ is an $X$-set of $G$, then $S\cap V(G_{i})$ is an $X$-set of $G_{i}$ for $i=1,\dots,k$.
While the focus of this section is on parameters defined to be the minimum cardinality over sets satisfying a given property, we begin with the more general Definition 2.2. This definition will facilitate an extension of the $X$-set idea to parameters defined to be the maximum cardinality over sets satisfying a given property; these are called $Y$-sets and are discussed in Section 8.
Definition 2.2.
A vertex set property $W$ is a property that is defined on subsets of the vertex set of each graph such that if $\varphi:V(G)\to V(G^{\prime})$ is a graph isomorphism, then for every $S\subseteq V(G)$ with property $W$, $\varphi(S)$ has property $W$. Let $W$ be a vertex set property and let $G$ be a graph. Then $S\subseteq V(G)$ is called an $W$-set if it has property $W$.
A vertex set property $W$ is called cohesive, or is a cohesive property, if every graph has at least one $W$-set. A graph parameter for which the value associated to each graph is solely determined by a cohesive property is called a cohesive parameter.
There are many natural graph parameters that can be associated to a cohesive property. For example, the minimum or maximum cardinality of an $W$-set, or the total number of $W$-sets. For a cohesive parameter, we can define the TAR graph of a base graph.
Definition 2.3.
For a cohesive $W$-set property $W$, the token addition and removal reconfiguration graph (TAR graph) of a base graph $G$ is the graph defined as follows: The vertex set of the TAR graph is the set of all $W$-sets of $G$. There is an edge between two vertices $S_{1}$ and $S_{2}$ of the TAR graph of $G$ if and only if $|S_{1}\ominus S_{2}|=1$.
The Superset axiom is the most fundamental of the original $X$-set axioms (for parameters that take the minimum cardinality of an $X$-set) and is needed for almost every result established in [6, 7]. For example, the Superset axiom ensures that the TAR reconfiguration graph is connected, making the study of connectedness of certain subgraphs of the TAR graph meaningful. So we give a name to a cohesive parameter that is the minimum size of an $X$-set and satisfies the Superset axiom (and no other assumptions). Throughout what follows, we will always use an adjective to modify $X$-set parameter, such as original, super, or the name of one of the new variants (all of which are super $X$-set parameters). However, we will use the same notation for the TAR graph and additional parameters derived from any type of $X$-set parameter. Since the only notation used for reconfiguration graphs is for TAR graphs, we have omitted the superscript TAR used in [6, 7].
Definition 2.4.
A super $X$-set parameter is a cohesive parameter $X$ such that $X(G)$ is defined to be the minimum cardinality of an $X$-set of $G$ where the $X$-sets of $G$ satisfy the following condition:
  1. (I)
    (Superset) If $S$ is an $X$-set and $S\subset S^{\prime}$, then $S^{\prime}$ is an $X$-set.
When $X$ is a super $X$-set parameter, the $X$-TAR graph of a base graph $G$ is denoted by $\mathfrak{X}(G)$, and the upper $X$ number, denoted by $\overline{X}(G)$, is the maximum cardinality of a minimal $X$-set.
The next set of axioms allows recovery of all main $X$-set parameter results in [6, 7], and all these axioms are needed for many of the main results in full generality. Removal of the Empty set axiom does not require changes to most of the proofs, and allows application to skew forcing. Removal of the Isolated vertex axiom (which states that $X(K_{1})=1$) requires more changes but also allows the main results to be generalized to a larger class of parameters, including vertex cover number. Although the inclusion of the Superset axiom in the next definition is redundant, we think it is preferable to explicitly list all required axioms.
Definition 2.5.
A robust $X$-set parameter is a super $X$-set parameter $X$ such that $X(G)$ and the $X$-sets of $G$ satisfy the following conditions:
  1. (I)
    (Superset) If $S$ is an $X$-set of $G$ and $S\subseteq S^{\prime}$, then $S^{\prime}$ is an $X$-set of $G$.
  2. (II)
    ($(n-1)$-set) If $G$ is a connected graph of order $n\geq 2$, then every set of $n-1$ vertices is an $X$-set.
  3. (III)
    (Component consistency) Suppose $G=G_{1}\,\sqcup\,\dots\,\sqcup\,G_{k}$ where $G_{i},i=1,\dots,k$ are the connected components of $G$. Then $S$ is an $X$-set of $G$ if and only if $S\cap V(G_{i})$ is an $X$-set of $G_{i}$ for $i=1,\dots,k$.
Observation 2.6.
Let $X$ be a robust $X$-set parameter. Note that if $G$ has no isolated vertices, then every set of $|V(G)|-1$ vertices is an $X$-set.
From the previous observation, we see that the definition of robust $X$-set is equivalent to that obtained by replacing axiom (II) of Definition 2.5 by (ii) from Definition 2.1 (and keeping the other two axioms of Definition 2.5 unchanged). The change in the $(n-1)$-set axiom was made because the behavior of $X$-sets is different when $X(K_{1})=0$ compared to that of an original $X$-set parameter (which implies $X(K_{1})=1$). Both cases $X(K_{1})=1$ and $X(K_{1})=0$ are discussed in the next remark.
Remark 2.7.
Let $X$ be a robust $X$-set parameter. Then one of two things happens regarding isolated vertices for every graph $G$:
  1. $(1)$
    [6, Remark 2.3] $X(K_{1})=1$: Every isolated vertex of $G$ is in every $X$-set of $G$. In this case, if $G^{\prime}=G\sqcup rK_{1}$ (where $G$ has no isolated vertices), then $X(G^{\prime})=X(G)+r$ and $\mathfrak{X}(G^{\prime})\cong\mathfrak{X}(G)$.
  2. $(2)$
    $X(K_{1})=0$: No isolated vertex of $G$ is in any minimal $X$-set of $G$. Every set of $|V(G)|-1$ vertices is an $X$-set (even if $G$ has isolated vertices). Suppose $G^{\prime}=G\sqcup rK_{1}$ (where $G$ has no isolated vertices). Then $X(G^{\prime})=X(G)$. However, $\mathfrak{X}(K_{1})\cong K_{2}$ when $X(K_{1})=0$. We see in the next proposition that $\mathfrak{X}(G^{\prime})\cong\mathfrak{X}(G)\Box K_{2}\Box\cdots\Box K_{2}$ ($r$ copies of $K_{2}$).
Therefore, it is sufficient to study TAR reconfiguration graphs of base graphs with no isolated vertices for robust $X$-set parameters.
Case (1) of the previous remark applies to all parameters studied in [6] and [7]. The vertex cover number is an example of a robust $X$-set parameter where case (2) applies.
The next result was established for original $X$-set parameters in Proposition 1.4 of [7]. It reduces consideration of a TAR reconfiguration problem to connected base graphs.
Proposition 2.8.
Let $X$ be a super $X$-set parameter that satisfies the Component consistency axiom and let $G=G_{1}\,\sqcup\,G_{2}$. Then $\mathfrak{X}(G)\,\cong\,\mathfrak{X}(G_{1})\,\Box\,\mathfrak{X}(G_{2})$.
Proof.
Let $S$ and $T$ be $X$-sets of $G$. Then $S=S_{1}\sqcup S_{2}$ and $T=T_{1}\sqcup T_{2}$, where $S_{1}$ and $T_{1}$ are $X$-sets of $G_{1}$, and $S_{2}$ and $T_{2}$ are $X$-sets of $G_{2}$. Observe that $S$ and $T$ are adjacent in $\mathfrak{X}(G)$ if and only if there exists a vertex $v\in V(G_{1})\sqcup V(G_{2})$ such that
  • $S_{1}=T_{1}$ and ($T_{2}=S_{2}\setminus\{v\}$ or $S_{2}=T_{2}\setminus\{v\}$), or
  • $S_{2}=T_{2}$ and ($T_{1}=S_{1}\setminus\{v\}$ or $S_{1}=T_{1}\setminus\{v\}$).
Further, $T_{2}=S_{2}\setminus\{v\}$ or $S_{2}=T_{2}\setminus\{v\}$ if and only if $S_{2}$ is adjacent to $T_{2}$ in $\mathfrak{X}(G_{2})$, and $T_{1}=S_{1}\setminus\{v\}$ or $S_{1}=T_{1}\setminus\{v\}$ if and only $S_{1}$ is adjacent to $T_{1}$ in $\mathfrak{X}(G_{1})$. ∎
As noted earlier, every original $X$-set parameter is a robust $X$-set parameter. This includes (standard) zero forcing, PSD forcing, domination, and power domination, as noted in [7]. Skew forcing is a robust $X$-set parameter that is not an original $X$-set parameter. The vertex cover number is another (very different) robust $X$-set parameter.
Of course, one can also define examples of robust $X$-set parameters that have no known use, as in the next example.
Example 2.9.
Given a connected graph $G$ of order $n\geq 2$, a set $S\subseteq V(G)$ is a star-set of $G$ if and only if $|S|\geq|V(G)|-1$. The only star-set of $K_{1}$ is its vertex. Given a disconnected graph $G$, a set $S\subseteq V(G)$ is a star-set of $G$ if and only if $S\cap V(C)$ is a star-set of $C$ for every connected component $C$ of $G$. Define $X^{*}(G)$ to be the minimum cardinality of a star-set of $G$. It is clear that for any graph $G$ of order $n$, $X^{*}(G)=n-c$ where $c$ is the number of connected components of $G$ of order at least two. Furthermore, for every connected graph of order $n$, the star reconfiguration graph is $\mathfrak{X}^{*}(G)\cong K_{1,n}$.
It is worth noting that the Component consistency axiom is irrelevant for connected graphs, which leads to the next definition and observation.
Definition 2.10.
A connected $X$-set parameter is a super $X$-set parameter $X$ such that $X(G)$ and the $X$-sets of $G$ satisfy the following conditions:
  1. (I)
    (Superset) If $S$ is an $X$-set of $G$ and $S\subseteq S^{\prime}$, then $S^{\prime}$ is an $X$-set of $G$.
  2. (II)
    (($n-1$)-sets) If $G$ is a connected graph of order $n\geq 2$, then every set of $n-1$ vertices is an $X$-set.
Observation 2.11.
If $X$ is a connected graph parameter, then every result that is true for a robust $X$-set parameter and graphs with no isolated vertices holds for $X$ for connected graphs of order at least two.
While most of the main results of this section are stated for robust parameters and graphs with no isolated vertices, they remain true for connected graphs and connected parameters as noted in Observation 2.11.
2.1 Initial Results
Most of the results in this section were established in [6] for original $X$-set parameters. The first statement in the next remark is Observation 2.4 of [6].
Remark 2.12.
Let $X$ be a super $X$-set parameter. Knowledge of all the minimal $X$-sets is sufficient to determine $\mathfrak{X}(G)$.
Suppose $G$ is a graph and $X_{1}$ and $X_{2}$ are super $X$-set parameters such that $M\subseteq V(G)$ is a minimal $X_{1}$-set of $G$ if and only if minimal $X_{2}$-set of $G$. Since the TAR graph is determined by the minimal $X$ sets, $\mathfrak{X}_{1}(G)=\mathfrak{X}_{2}(G)$.
Since several of the parameters we study in the following sections have the same minimal $X$-sets for certain graph families, this allows the transfer of results about TAR graphs from one parameter to another for such families of graphs.
Remark 2.13.
Suppose $X$ is a super $X$-set parameter that also satisfies the $(n-1)$-set axiom and let $G$ be a graph of order at least two. If $X(K_{1})=0$ or if $G$ has no isolated vertices, then no vertex is in every minimal $X$-set. If in addition the empty set is not an $X$-set of $G$, then $G$ has more than one minimal $X$-set.
The inequality $\Delta(\mathfrak{X}(G))\leq n$ is true for all super $X$-set parameters and graphs, without additional assumptions, as seen in the proof of Proposition 2.5(1) in [6]. The next result also shows how closely related the $(n-1)$-set axiom is to $\Delta(G)=n$.
Proposition 2.14.
Let $X$ be a super $X$-set parameter and $G$ be a graph of order $n$. Then $\Delta(\mathfrak{X}(G))=n$ if and only if every set of $n-1$ vertices is an $X$-set.
Proof.
Let $S\subseteq V(G)$ be a vertex of $\mathfrak{X}(G)$ and let $|S|=r$. Then there are at most $r$ vertices that can be deleted (one at a time) leaving an $X$-set, and at most $n-r$ vertices that can be added (one at a time). Thus $\deg_{\mathfrak{X}(G)}(S)\leq n$. It is immediate that $\Delta(\mathfrak{X}(G))=n$ if every set of $n-1$ vertices is an $X$-set.
Suppose that $\Delta(\mathfrak{X}(G))=n$. Then there exists an $X$-set $S$ such that $S\cup\{v\}$ and $S\setminus\{w\}$ are $X$-sets for every $v\notin S$ and $w\in S$. For $u\in V(G)$, define $R_{u}=V(G)\setminus\{u\}$. Then $R_{u}$ is a superset of $S$ if $u\notin S$ and is a superset of $S\setminus\{u\}$ if $u\in S$. In either case, $R_{u}$ is an $X$-set by the Superset axiom. ∎
The next result is central to many of the results that follow, including the isomorphism results. If $X(K_{1})=0$ or $G$ has no isolated vertices, then every set of $n-1$ vertices is an $X$-set by Remark 2.7 or Observation 2.6, so $\deg_{\mathfrak{X}(G)}(V(G))=n$.
Corollary 2.15.
Let $X$ be a robust $X$-set parameter. If $X(K_{1})=0$ or $G$ is a graph of order $n$ with no isolated vertices, then $\Delta(\mathfrak{X}(G))=n.$
Example 2.16 illustrates the necessity of the $(n-1)$-set axiom for the previous result and for results that depend on it.
Example 2.16.
Define $X_{V}$ to be the property that for every graph $G$, $V(G)$ is the one and only $X_{V}$-set, and define $X_{V}(G)$ to be the minimum cardinality of an $X_{V}$-set. Observe that $X_{V}$ is a super $X$-set parameter that stasfies all of the axioms for an original $X$-set parameter except the $(n-1)$-set axiom. Since $X_{V}(G)=|V(G)|$ and $\mathfrak{X}_{V}(G)=K_{1}$ for every graph $G$, for a graph $G$ with no isolated vertices, neither the order of $G$ nor $X_{V}(G)$ can be determined from $\mathfrak{X}_{V}(G)$. Many subsequent results also fail for $X_{V}$.
The connected domination number number, which is discussed in Section 7, provides an example of a parameter in the literature that is a super $X$-set parameter but not a robust $X$-set parameter, and for which Corollary 2.15 and TAR graph isomorphism results fail.
The next result was established for original $X$-set parameters in Proposition 2.6 of [6] (and used in the proof of what is Proposition 2.31 here). Since only the last statement was explicitly stated and proved, we included the brief revised proof here.
Proposition 2.17.
Let $X$ be a super $X$-set parameter, let $G$ be a graph on $n$ vertices, and let $S$ be an $X$ set. Then $\deg_{\mathfrak{X}(G)}(S)\geq n-|S|$ and $S$ is minimal if and only if $\deg_{\mathfrak{X}(G)}(S)=n-|S|$. Furthermore, $\delta(\mathfrak{X}(G))=n-\overline{X}(G)$.
Proof.
Let $\mathfrak{X}=\mathfrak{X}(G)$. Let $M$ be a minimal $X$-set of $G$, with $|M|=m$ for some $m\geq 0$. Since $N_{\mathfrak{X}}(M)=\{M\cup\{v\}:v\in V(G)\setminus M\}$, $n-m=\deg_{\mathfrak{X}}(M)$.
Suppose $S$ is not minimal and choose a minimal $X$-set $M\subsetneq S$. Denote the elements of $S\setminus M$ by $U=\{u_{1},\ldots,u_{\ell}\}$ and those of $V(G)\setminus S$ by $W=\{w_{1},\ldots,w_{n-m-\ell}\}$. Then the sets of the form $S\setminus\{u_{i}\}$ with $u_{i}\in U$ and $S\cup\{w_{j}\}$ with $w_{j}\in W$ are neighbors of $S$ in ${\mathfrak{X}}$, and $|N_{\mathfrak{X}}(S)|\geq\ell>n-m$. The last statement is now immediate. ∎
Additional minor results adapted from [6] can be found in Section 2.5.
2.2 Hypercube representation and applications
For $d\geq 1$, the graph having as vertices all $d$-tuples with entries in $\{0,1\}$ with two such sequences adjacent if and only if they differ in exactly one place is a characterization of $Q_{d}$, the hypercube of dimension $d$ (and $Q_{0}\cong K_{1}$). There is a well-known representation of any TAR reconfiguration graph as a subgraph of a hypercube. Let $G$ be a graph with $V(G)=\{v_{1},\dots,v_{n}\}$. Any subset $S$ of $V(G)$ can be represented by a sequence $(s_{1},\dots,s_{n})$ where $s_{i}=1$ if $v_{i}\in S$ and $s_{i}=0$ if $v_{i}\not\in S$. The first part of the next remark appeared in Remark 2.12 of [6]
Remark 2.18.
Let $X$ be a super $X$-set parameter. For a graph $G$ of order $n$, $\mathfrak{X}(G)$ is isomorphic to a subgraph of $Q_{n}$, and thus $\mathfrak{X}(G)$ is bipartite. If $G$ is a graph such that any one vertex is an $X$-set and the emptyset is not an $X$-set, then $\mathfrak{X}(G)\cong Q_{n}-v$ for any $v\in V(Q_{n})$.
Remark 2.19.
It is asserted in Remark 2.13 of [6] that the $X$-TAR graph is not a hypercube for an original $X$-set parameter and base graph of order at least two. This is not technically true: If $X$ is an original $X$-set parameter, then $\mathfrak{X}(nK_{1})\cong Q_{0}$. However, the intended statement, the $X$-TAR graph is not a hypercube for an original $X$-set parameter and base graph that has an edge, is true and the argument in [6, Remark 2.13] justifies that statement.
However, a robust $X$-set parameter such as skew forcing, may allow the $X$-TAR graph of a base graph with an edge to be a hypercube: Any graph $G$ such that $\emptyset$ is an $X$-set has $\mathfrak{X}(G)\cong Q_{n}$ for a robust $X$-set parameter. The $(n-1)$-set axiom also plays a role, since the super $X$-set parameter $X_{V}$ in Example 2.16 satisfies the (original) Isolated vertex, Connected component, and Empty set axioms but not the $(n-1)$-set axiom, and $\mathfrak{X}_{V}(G)\cong Q_{0}$ for all graphs $G$.
The next lemma was established for original $X$-set parameters in Lemma 2.14 of [6]; its corollary is immediate and appears for original $X$-set parameters in Corollary 2.15 of [6].
Lemma 2.20.
Let $X$ be a super $X$-set parameter and let $G$ be a graph on $n$ vertices and let $t\leq n$. Then $X(G)\leq n-t$ if and only if $\mathfrak{X}(G)$ has an induced subgraph isomorphic to the hypercube $Q_{t}$.
Proof.
Let $S$ be a minimum $X$-set and let $W=V(G)\setminus S$. The induced subgraph of $\mathfrak{X}(G)$ having vertices consisting of sets of the form $S\sqcup W^{\prime}$ over all subsets $W^{\prime}\subseteq W$ is $Q_{n-X(G)}$. Any hypercube $Q_{t}$ for $n-X(G)\geq t\geq 0$ is an induced subgraph of $Q_{n-X(G)}$.
Suppose $H$ is an induced subgraph of $\mathfrak{X}(G)$ isomorphic to $Q_{t}$ for some $1\leq t\leq n$. Choose $S\in V(H)$ such that $|S|$ is minimum over all $X$-sets in $H$. Since no vertex in $H$ has fewer vertices than $S$, every one of the $t$ neighbors of $S$ in $H$ is obtained by adding a vertex of $G$ to $S$. Thus $|V(G)\setminus S|\geq t$ and $X(G)\leq|S|\leq n-t$. ∎
Corollary 2.21.
Let $X$ be a super $X$-set parameter and let $G$ be a graph on $n$ vertices. Then $d=n-X(G)$ is the maximum dimension of a hypercube isomorphic to an induced subgraph of the reconfiguration graph $\mathfrak{X}(G)$.
The next result follows from prior results and was established for original $X$-set parameters in Corollary 2.16 of [6].
Corollary 2.22.
Let $X$ be a robust $X$-set parameter and let $G$ and $G^{\prime}$ be graphs such that $\mathfrak{X}(G)\cong\mathfrak{X}(G^{\prime})$. Suppose further that $X(K_{1})=0$ or $G$ and $G^{\prime}$ have no isolated vertices. Then
  1. 1.
    $|V(G)|=|V(G^{\prime})|$.
  2. 2.
    $X(G)=X(G^{\prime})$.
  3. 3.
    $\overline{X}(G)=\overline{X}(G^{\prime})$.
Proof.
By Corollary 2.15, $|V(G)|=\Delta(\mathfrak{X}(G))=\Delta(\mathfrak{X}(G^{\prime}))=|V(G^{\prime})|$. Corollary 2.21 implies $X(G)=X(G^{\prime})$. By Proposition 2.17, $\overline{X}(G)=n-\delta(\mathfrak{X}(G))=n-\delta(\mathfrak{X}(G^{\prime}))=\overline{X}(G^{\prime})$. ∎
2.3 Isomorphism results
The original $X$-set parameter version of the next result is Theorem 1.1 of [7], and we have verified that the expanded version including $X(K_{1})=0$ remains true for robust $X$-set parameters. The proof is established through a series of preliminary results, which are stated for robust $X$-set parameters (or super $X$-set satisfying some additional axioms) after Theorem 2.23.
Theorem 2.23.
Let $X$ be a robust $X$-set parameter and let $G$ and $G^{\prime}$ be base graphs such that $\mathfrak{X}(G)\cong\mathfrak{X}(G^{\prime})$. If $X(K_{1})=0$ or $G$ and $G^{\prime}$ have no isolated vertices, then there is a relabeling of the vertices of $G^{\prime}$ such that $G$ and $G^{\prime}$ have exactly the same $X$-sets.
Observe that Proposition 2.14 shows that the $(n-1)$-set axiom (or at least something beyond the Superset axiom) is necessary for the previous result, as well as Theorem 2.27 (which together with Theorem 2.28 and Remark 2.29 implies Theorem 2.23).
Next we reproduce some necessary definitions from [7]. Let $P=(Y,\leq)$ be a poset. An element $u\in Y$ is maximal (respectively, minimal) if $u\not<y$ (respectively, $u\not>y$) for each $y\in Y$. Two elements $u$ and $v$ of $Y$ are comparable provided $u\leq v$ or $v\leq u$. A poset such that each pair of elements is comparable is called a chain. An interval in $P$ is a set $[u,v]=\{y:u\leq y\leq v\}$ where $u\leq v$. The length of a chain $C$ is $|C|-1$. The length of an interval $[u,v]$ is the maximum length of a chain in $[u,v]$ and is denoted by $\ell(u,v)$. For a set $T$, the power set of $T$ is denoted by $\mathcal{P}(T)$. If $X$ is a super $X$-set property, then $\mathfrak{X}(G)$ is a join-semilattice contained in $(\mathcal{P}(V(G)),\subseteq)$ and $\ell(S,S^{\prime})=\operatorname{dist}_{\mathfrak{X}(G)}(S,S^{\prime})$ for any interval $[S,S^{\prime}]$ in $\mathfrak{X}(G)$ since $S$ and $S^{\prime}$ are comparable. The next results was established for original $X$-set parameters in Lemma 2.1 of [7].
Lemma 2.24.
Let $X$ be a super $X$-set parameter, let $t\geq 0$ be an integer, let $G$ be a graph on $n$ vertices, and let $H$ be an induced subgraph of $\mathfrak{X}(G)$. If $H\cong Q_{t}$, then $V(H)$ is an interval of length $t$ in the poset $(\mathcal{P}(V(G)),\subseteq)$.
Proof.
Assume $H\cong Q_{t}$. The claim is obvious for $t=0,1$, so suppose that $t\geq 2$. We begin by showing that $(V(H),\subseteq)$ has exactly 1 maximal element. Assume, to the contrary, that $(V(H),\subseteq)$ has at least 2 maximal elements $u$ and $v$. Since $H$ is connected, there exists a path from $u$ to $v$ in $H$. Every path $P$ from $u$ to $v$ in $H$ can be written in the form $(u,y_{1},\dots,y_{i},v)$. For each such $P$, let $d_{P}$ to be the smallest index $k$ such that $y_{k-1}\subseteq y_{k}$, where $y_{0}=u$ and $y_{i+1}=v$. Note that since $u$ and $v$ are maximal in $(V(H),\subseteq)$, $d_{P}\geq 2$ is always defined.
Let $d$ be the minimum $d_{P}$ amongst all paths $P$ from $u$ to $v$ in $H$. Pick a path $P^{*}$ in $H$ of the form $(u,x_{1},\dots,x_{i},v)$ such that $d_{P^{*}}=d$, and let $x_{0}=u$ and $x_{i+1}=v$. Then $x_{d-1}\subseteq x_{d-2}$ because $x_{d-2}\not\subseteq x_{d-1}$ and $x_{d-2}$ is adjacent to $x_{d-1}$. Since $x_{d-2}$ and $x_{d}$ have $x_{d-1}$ as a common neighbor and each pair of vertices in a hypercube share exactly 0 or 2 common neighbors, there exists some vertex $w\not=x_{d-1}$ in $H$ that is adjacent to $x_{d-2}$ and $x_{d}$. By our choice of $d$, $w\subseteq x_{d-2}$ and $w\subseteq x_{d}$ or else $P^{\prime}=(u,x_{1},\dots,x_{d-2},w,x_{d},\dots,x_{i},v)$ would be a path with $d_{P^{\prime}}<d_{P^{*}}$ since $w=x_{d-2}\cup x_{d}$. Hence $x_{d-2}\cap x_{d}=w$. This is absurd since $x_{d-2}\cap x_{d}=x_{d-1}$ and $w\not=x_{d-1}$.
Thus, $H$ has exactly 1 maximal element $T$. A similar argument shows $H$ has exactly 1 minimal element $S$ in the subset partial ordering of $\mathfrak{X}(G)$. Thus, $S\subseteq R\subseteq T$ for every $R\in V(H)$. Since $H\cong Q_{t}$, $\operatorname{dist}(S,T)\leq\operatorname{diam}(Q_{t})=t$. So there are at most $2^{t}$ elements in the interval $[S,T]$. But $|V(H)|=|V(Q_{t})|=2^{t}$. Therefore, $\operatorname{dist}(S,T)=t$ and $V(H)=[S,T]$.∎
The proof of Lemma 2.2 in [7] also establishes the next result.
Lemma 2.25.
Let $X$ be a super $X$-set parameter and let $G$ and $G^{\prime}$ be graphs. Suppose $\varphi:\mathfrak{X}(G)\to\mathfrak{X}(G^{\prime})$ is a graph isomorphism. Let $S^{\prime}=\varphi(V(G))$. If $M^{\prime}$ is a minimal $X$-set of $G^{\prime}$, then $M^{\prime}\subseteq S^{\prime}$.
Proof.
Let $M^{\prime}$ be a minimal $X$-set of $G^{\prime}$. Define $M=\varphi^{-1}(M^{\prime})$. The interval $[M,V(G)]$ in $(\mathcal{P}(V(G),\subseteq)$ forms an induced $Q_{t}$ in $\mathfrak{X}(G)$ for some integer $t>0$. By Lemma 2.24, $\varphi([M,V(G)])$ is an interval $[Z^{\prime},W^{\prime}]$ in $(\mathcal{P}(V(G^{\prime})),\subseteq)$. Since $M^{\prime}$ is a minimal $X$-set of $G^{\prime}$, $Z^{\prime}=M^{\prime}$. Thus, $M^{\prime}\subseteq S^{\prime}$. ∎
Irrelevant vertices play a key role in the proofs of results about TAR-graph isomorphsims of original $X$-set parameters in [7]. Here we extend the definition of irrelevant set to super $X$-set parameters and reproduce results that remain true (some results require the parameter to be robust). Let $G$ be a graph and let $X$ be a super $X$-set parameter. A vertex $v\in V(G)$ is $X$-irrelevant if $v\not\in S$ for every minimal $X$-set $S$ of $G$. Observe that if $X(K_{1})=0$ and $X$ satisfies the Component consistency axiom, then every isolated vertex is irrelevant. A set $R\subseteq V(G)$ is an $X$-irrelevant set if every vertex of $R$ is $X$-irrelevant. For a graph $G$ and $R\subseteq V(G)$, define the map $\nu_{R}:V(\mathfrak{X}(G))\to\mathcal{P}(V(G))$ by $\nu_{R}(S)=S\ominus R$. The proof of Theorem 2.7 in [7] for original $X$-set parameters also establishes the next result.
Theorem 2.26.
Let $X$ be a super $X$-set parameter, let $G$ be a graph, and let $R\subseteq V(G)$. Then $\nu_{R}$ is a graph automorphism of $\mathfrak{X}(G)$ if and only if $R$ is $X$-irrelevant.
Proof.
Suppose that $\nu_{R}$ is an automorphism of $\mathfrak{X}(G)$. By Lemma 2.25, every minimal $X$-set of $G$ is a subset of $\nu_{R}(V(G))=V(G)\setminus R$. Thus, $R$ is $X$-irrelevant.
Suppose that $R$ is $X$-irrelevant. Let $S$ be an $X$-set of $G$. Then there exists some minimal $X$-set $T\subseteq S$. Since $R$ is $X$-irrelevant, $T\subseteq S\setminus R\subseteq S$. Thus, $\nu_{R}(S)\supseteq S\setminus R\supseteq T$ is an $X$-set of $G$. Adjacency in $\mathfrak{X}(G)$ is preserved by $\nu_{R}$: Consider the adjacent $X$-sets $S$ and $S^{\prime}=S\cup\{u\}$. If $u\neq R$, then $S^{\prime}\setminus R=(S\setminus R)\cup\{u\}$, so $\nu_{R}(S^{\prime})$ is adjacent to $\nu_{R}(S)$. If $u\in R$, then $S\setminus R=(S^{\prime}\setminus R)\cup\{u\}$, so $\nu_{R}(S)$ is adjacent to $\nu_{R}(S^{\prime})$. Therefore, $\nu_{R}$ is an automorphism. ∎
Theorems 2.27 and 2.28 (and Remark 2.29) imply Theorem 2.23. The proof of Theorem 2.8 [7] also establishes Theorem 2.27 (even with the addition of $X(K_{1})=0$ case). We supply a proof for Theorem 2.28 that is modified from the proof of Theorem 2.9 in [7] to weaken the hypotheses so that $X$ need only be a super $X$-set parameter rather than a robust $X$-set parameter provided the two graphs have the same order.
Theorem 2.27.
Let $X$ be a robust $X$-set parameter, let $G$ and $G^{\prime}$ be graphs, and let $\tilde{\varphi}:\mathfrak{X}(G)\to\mathfrak{X}(G^{\prime})$ be an isomorphism. Suppose further that $X(K_{1})=0$ or $G$ and $G^{\prime}$ have no isolated vertices. Then $R^{\prime}=V(G^{\prime})\setminus\tilde{\varphi}(V(G))$ is $X$-irrelevant and $\varphi=\nu_{R^{\prime}}\circ\tilde{\varphi}$ is an isomorphism such that $|\varphi(S)|=|S|$ for every $S\in V(\mathfrak{X}(G))$.
Proof.
Let $n=|V(G)|$. If $\tilde{\varphi}(V(G))=V(G^{\prime})$, then $R^{\prime}=\emptyset$ is $X$-irrelevant. So suppose that $\tilde{\varphi}(V(G))=S^{\prime}$, where $S^{\prime}\not=V(G^{\prime})$. Since $\tilde{\varphi}$ is an isomorphism, Lemma 2.25 implies every minimal $X$-set of $G^{\prime}$ is a subset of $S^{\prime}$ and hence $R^{\prime}$ is $X$-irrelevant. By Theorem 2.26, $\nu_{R^{\prime}}$ is an automorphism of $\mathfrak{X}(G^{\prime})$. Thus, $\varphi=\nu_{R^{\prime}}\circ\tilde{\varphi}$ is an isomorphism such that $\varphi(V(G))=V(G^{\prime})$.
Note that $|V(G^{\prime})|=n$ and $X(G^{\prime})=X(G)$ by Corollary 2.22. Let $S\in V(\mathfrak{X}(G))$. The interval $H=[S,V(G)]\in\mathfrak{X}(G)$ is an induced hypercube in $\mathfrak{X}(G)$, so $\varphi(H)$ is an induced hypercube in $\mathfrak{X}(G^{\prime})$. By Lemma 2.24 and since $\varphi(V(G))=V(G^{\prime})$, $\varphi(H)=[S^{\prime},V(G^{\prime})]$ in $\mathfrak{X}(G^{\prime})$ and $\operatorname{dist}(S^{\prime},V(G^{\prime}))=\operatorname{dist}(S,V(G))$. We show by induction on $|S|$ that $|\varphi(S)|=|S|$. We say $S$ is a $k$-$X$-set if $S$ is an $X$-set and $|S|=k$.
For the base case, assume $|S|=X(G)$, so $\operatorname{dist}(S^{\prime},V(G^{\prime}))=\operatorname{dist}(S,V(G))=n-X(G)=|V(G^{\prime})|-X(G^{\prime})$. This implies $|S^{\prime}|=X(G^{\prime})=X(G)=|S|$. The same reasoning applies using $\varphi^{-1}$, since $\varphi^{-1}(V(G^{\prime}))=V(G)$. Thus $\varphi$ defines a bijection between minimum $X$-sets of $G$ and minimum $X$-sets of $G^{\prime}$.
Now assume $\varphi$ defines a bijection between $i$-$X$-sets of $G$ and $i$-$X$-sets of $G^{\prime}$ for $X(G)\leq i\leq k$ and let $S$ be a $(k+1)$-$X$-set of $G$. This implies $|\varphi(W)|\geq k+1$ for $W\in[S,V(G)]$. By Lemma 2.24, $\varphi([S,V(G)])=[S^{\prime},V(G^{\prime})]$ in $\mathfrak{X}(G^{\prime})$ and $\operatorname{dist}(S^{\prime},V(G^{\prime}))=n-k-1$. Thus $S^{\prime}$ is a $(k+1)$-$X$-set of $G$. ∎
For any map $\psi:A\to A^{\prime}$ and subset $B\subseteq A$ we write $\psi(B)$ to mean the image of $B$, i.e., $\psi(B)=\{\psi(b):b\in B\}$. This is particularly useful when working with a map $\psi:V(G)\to V(G^{\prime})$ that maps $X$-sets of $G$ to $X$-sets of $G^{\prime}$, since this convention naturally induces a map $\psi:V(\mathfrak{X}(G))\to V(\mathfrak{X}(G^{\prime}))$.
Theorem 2.28.
Let $X$ be a super $X$-set parameter, let $G$ and $G^{\prime}$ be graphs of order $n$, and suppose $\varphi:\mathfrak{X}(G)\to\mathfrak{X}(G^{\prime})$ is a graph isomorphism.
Then $|\varphi(S)|=|S|$ for every $X$-set $S$ of $G$ if and only if there exists a bijection $\psi:V(G)\to V(G^{\prime})$ such that $\psi(S)=\varphi(S)$ for every $X$-set $S$ of $G$.
If $\varphi:\mathfrak{X}(G)\to\mathfrak{X}(G^{\prime})$ is a graph isomorphism such that $|\varphi(S)|=|S|$ for every $X$-set $S$, then there is a relabeling of the vertices of $G^{\prime}$ such that the relabeled graph has the same $X$-sets as $G$ and the same $X$-TAR graph as $G$.
Proof.
If there exists a bijection $\psi:V(G)\to V(G^{\prime})$ such that $\psi(S)=\varphi(S)$ for every $X$-set $S$ of $G$, then it is immediate that $|\varphi(S)|=|S|$ for every $X$-set $S$ of $G$. Assume $|\varphi(S)|=|S|$ for every $X$-set $S$ of $G$. For any graph $H$ and $v\in V(H)$, define $R_{v}=V(H)\setminus\{v\}$. If $R_{v}$ is an $X$-set of $G$, let $v^{\prime}$ be the unique vertex such that $\varphi(R_{v})=R_{v^{\prime}}$. Let $W$ be the set of all vertices $w\in V(G)$ such that $R_{w}$ is not an $X$-set of $G$ and let $W^{\prime}$ be the set of all vertices $w^{\prime}\in V(G^{\prime})$ such that $R_{w^{\prime}}$ is not an $X$-set of $G^{\prime}$. Since $|\varphi(S)|=|S|$ for every $X$-set $S$ of $G$, $|W|=|W^{\prime}|$ and so there exists a bijection $\theta:W\to W^{\prime}$. Let $\psi:V(G)\to V(G^{\prime})$ be given by
$\psi(v)=\begin{cases}v^{\prime}&\text{if $S_{v}$ is an $X$-set of $G$,}\\ \theta(v)&\text{otherwise.}\end{cases}$
Observe that $\psi:V(G)\to V(G^{\prime})$ is a bijection. The proof that $\psi(S)=\varphi(S)$ for every $X$-set $S$ of $G$ is as the proof of Theorem 2.9 in [7], proceeding iteratively from $|S|=n$ to $|S|=X(G)$.
By the choice of $\varphi$ and the definition of $\psi$, we have $\varphi(S)=\psi(S)$ for $|S|=n,n-1$. Assume $\varphi(S)=\psi(S)$ for each $X$-set $S$ of order $k$ for some $k$ with $n-1\geq k>X(G)$. Let $S$ be an $X$-set of order $k-1$. Since $k-1\leq n-2$, there exist distinct vertices $a,b\in V(G)\setminus S$ such that $S\cup\{a\}$ and $S\cup\{b\}$ are $X$-sets of order $k$. Since $S$ is adjacent to $S\cup\{a\}$ and $S\cup\{b\}$ in $\mathfrak{X}(G)$, and $|\varphi(S\cup\{a\})|=|\varphi(S\cup\{b\})|>|\varphi(S)|$, there exist distinct $a^{\prime},b^{\prime}\in V(G^{\prime})\setminus\varphi(S)$ such that $\varphi(S\cup\{a\})=\varphi(S)\cup\{a^{\prime}\}$ and $\varphi(S\cup\{b\})=\varphi(S)\cup\{b^{\prime}\}$. Thus,
$\varphi(S)=\varphi(S\cup\{a\})\cap\varphi(S\cup\{b\})=\psi(S\cup\{a\})\cap\psi(S\cup\{b\})=\psi(S),$
where the last equality follows since $\psi$ is a bijection.
Similarly, the proof that if $\varphi:\mathfrak{X}(G)\to\mathfrak{X}(G^{\prime})$ is a graph isomorphism such that $|\varphi(S)|=|S|$ for every $X$-set $S$, then there is a relabeling of the vertices of $G^{\prime}$ such that the relabeled graph has the same $X$-sets as $G$ and the same $X$-TAR graph as $G$ follows the proof of Theorem 2.9 in [7]. Finally, suppose $|\varphi(S)|=|S|$ for every $X$-set $S$ and $\psi:V(G)\to V(G^{\prime})$ is a bijection such that $\psi(S)=\varphi(S)$ for every $X$-set $S$ of $G$. Define $G^{\prime\prime}$ from $G^{\prime}$ by relabeling vertices of $G^{\prime}$ so that $v^{\prime}\in V(G^{\prime})$ is labeled by $\psi^{-1}(v^{\prime})$. Then $G^{\prime\prime}\cong G^{\prime}$ and $G^{\prime\prime}$ and $G$ have the same $X$-sets.
Remark 2.29.
Let $X$ be a robust $X$-set parameter, and let $G$ and $G^{\prime}$ be graphs such that $\varphi:\mathfrak{X}(G)\to\mathfrak{X}(G^{\prime})$ is a graph isomorphism. If $X(K_{1})=0$ or $G$ and $G^{\prime}$ have no isolated vertices, then Corollary 2.22 implies that the hypotheses of Theorem 2.28 are satisfied.
2.4 $X$-irrelevant vertices and automorphisms of $X$-TAR graphs.
In this section we point out that the proof of the characterization of automorphisms of $X$-TAR graphs in Theorem 2.13 in [7] remains true for robust parameters. Let $M_{X}(G)$ denote the set of bijections $\psi:V(G)\to V(G)$ that send minimal $X$-sets of $G$ to minimal $X$-sets of $G$ of the same size.
Theorem 2.30.
Let $X$ be a robust $X$-set parameter and let $G$ be a graph. If $X(K_{1})=0$ or $G$ has no isolated vertices, then the automorphism group of $\mathfrak{X}(G)$ is generated by
$\{\nu_{R}:R\text{ is $X$-irrelevant}\}\cup M_{X}(G).$
The proof of the previous result needs two preliminary results, The next proposition follows immediately from Proposition 2.17, which implies that whether or not a set is minimal can be determined from its cardinality and its degree; it is also stated and proved for $X$-sets as Proposition 2.10 in [7].
Proposition 2.31.
Let $X$ be a super $X$-set parameter, and let $G$ and $G^{\prime}$ be graphs. Suppose $\varphi:\mathfrak{X}(G)\to\mathfrak{X}(G^{\prime})$ is a graph isomorphism such that $|\varphi(S)|=|S|$. Then $\varphi$ maps minimal $X$-sets to minimal $X$-sets (of the same size).
The next result is established by the proof of Proposition 2.11 in [7] .
Proposition 2.32.
Let $X$ be a super $X$-set parameter, and let $G$ and $G^{\prime}$ be graphs. Suppose $\psi:V(G)\to V(G^{\prime})$ is a bijection.
  1. $(1)$
    Suppose $\psi$ maps $X$-sets of $G$ to $X$-sets of $G^{\prime}$. Then the induced mapping $\psi:V(\mathfrak{X}(G))\to V(\mathfrak{X}(G^{\prime}))$ is an isomorphism of $\mathfrak{X}(G)$ and $\psi(\mathfrak{X}(G))$. If every $X$-set of $G^{\prime}$ is the image of an $X$-set of $G$, then $\psi:V(\mathfrak{X}(G))\to V(\mathfrak{X}(G^{\prime}))$ is an isomorphism of $\mathfrak{X}(G)$ and $\mathfrak{X}(G^{\prime})$.
  2. $(2)$
    Suppose $\psi$ maps minimal $X$-sets of $G$ to minimal $X$-sets of $G^{\prime}$. Then $\psi$ maps $X$-sets of $G$ to $X$-sets of $G^{\prime}$. If every minimal $X$-set of $G^{\prime}$ is the image of a minimal $X$-set of $G$, then $\psi$ is a bijection from $X$-sets of $G$ to $X$-sets of $G^{\prime}$.
  3. $(3)$
[…]
Proof.
(1): Since $\psi$ maps $X$-sets of $G$ to $X$-sets of $G^{\prime}$, $\psi$ induces a bijection between the vertices of $\mathfrak{X}(G)$ and a subset of the vertices of $\mathfrak{X}(G^{\prime})$ ($X$-sets of $G^{\prime}$ of the form $\psi(S)$ where $S$ is an $X$-set of $G$). Assume that $S_{1},S_{2}\in V(\mathfrak{X}(G))$ are adjacent in $\mathfrak{X}(G)$. Without loss of generality, $|S_{1}\setminus S_{2}|=1$. Since $\psi$ is a bijection, $|\psi(S_{1})\setminus\psi(S_{2})|=1$. Thus, $\psi(S_{1})$ and $\psi(S_{2})$ are adjacent in $\mathfrak{X}(G^{\prime})$. Hence $\psi$ is an isomorphism from $\mathfrak{X}(G)$ to $\psi(\mathfrak{X}(G))$.
(2): Let $S\in V(\mathfrak{X}(G))$ be an $X$-set. There is a minimal $X$-set $T\subseteq S$ of $G$ and $\psi(T)\subseteq\psi(S)$. Since $\psi(T)$ is a minimal $X$-set of $G^{\prime}$, $\psi(S)$ is an $X$-set of $G^{\prime}$.
Statement (3) is immediate from statements (2) and (1) ∎
Proof of Theorem 2.30:
Proof.
By Theorem 2.26 $\nu_{R}\in\text{aut}(\mathfrak{X}(G))$ for every $X$-irrelevant set $R$. By Proposition 2.32(3), $\psi\in\text{aut}(\mathfrak{X}(G))$ for every $\psi\in M_{X}(G)$ (since $G^{\prime}=G$ here, having $\psi$ map minimal $X$-sets to minimal $X$-sets is sufficient).
We now show that $\{\nu_{R}:R\text{ is $X$-irrelevant}\}\cup M_{X}(G)$ generates aut$(\mathfrak{X}(G))$. Let $\varphi$ by an automorphism of $\mathfrak{X}(G)$. Suppose first that $V(G)$ is fixed by $\varphi$. Then $|S|=n-\operatorname{dist}(V(G),S)=n-\operatorname{dist}(V(G),\varphi(S))=|\varphi(S)|$ for each $S\in V(\mathfrak{X}(G))$. By Theorem 2.28 there exists a bijection $\psi:V(G)\to V(G)$ such that $\psi(S)=\varphi(S)$ for every $X$-set $S$ of $G$. By Proposition 2.31, $\psi$ maps minimal $X$-sets to minimal $X$-sets of the same size. Thus $\psi\in M_{X}$.
Suppose that $V(G)$ is not fixed by $\varphi$. By Theorem 2.27 and the preceding argument, there exists a bijection $\psi\in M_{X}(G)$ such that $\psi=\nu_{R}\circ\varphi$, where $R=V(G)\setminus\varphi(G)$. Thus, $\varphi=\nu_{R}^{-1}\circ\psi$. ∎
Recall that if $X(K_{1})=0$, then every isolated vertex is irrelevant (see Remark 2.7)
2.5 Connectedness
A main question in reconfiguration is: For which $k$ is the subgraph of the TAR reconfiguration graph induced by sets of no more than $k$ vertices connected? For a super $X$-set parameter $X$, $\mathfrak{X}(G)$ is always connected (every $X$-set can be augmented one vertex at a time to get $V(G)$). Parameters relating to connectedness were defined for domination and bounds on these parameters were established in [19]. The definitions and bounds were extended to original $X$-sets in [6]. In this section we further extend these definitions and results to super $X$-set parameters, showing they remain valid for a larger class of parameters that the isomorphism results. For example, these results apply to connected domination, which is not robust (see Section 7). A method involving twins was used in [7] to construct examples with strict inequality in one of the bounds, and we extend this from zero forcing to super $X$-set parameters that satisfy additional conditions on twins.
Definition 2.33.
Suppose $X$ is a super $X$-set parameter. The $k$-token addition and removal (TAR) reconfiguration graph for $X$, denoted by $\mathfrak{X}_{k}(G)$, is the subgraph of $\mathfrak{X}(G)$ induced by the set of all $X$-sets of cardinality at most $k$ as its vertex set.
The least $k_{0}$ such that $\mathfrak{X}_{k}(G)$ is connected for all $k\geq k_{0}$ is denoted by $x_{0}(G)$, and the least $k$ such that $\mathfrak{X}_{k}(G)$ is connected is denoted by $\underline{x_{0}}(G)$.
Knowing $x_{0}(G)$ or $\underline{x_{0}}(G)$ allows us to work within a smaller TAR reconfiguration graph when modifying one solution to another. Since $\mathfrak{X}_{k}(G)$ is an induced subgraph of $\mathfrak{X}(G)$, many elementary results concerning $\mathfrak{X}(G)$ also apply to $\mathfrak{X}_{k}(G)$
The next result was established for original $X$-set parameters in Proposition 2.8 in [6]. Although that result does not include Proposition 2.34(2), that statement is immediate since adding one vertex at a time to the one minimal $X$-set does not disconnect the graph. The proof of [6, Proposition 2.8] remains valid for any super $X$-set parameter (and the hypothesis that $G$ has no isolated vertices is not needed).
Proposition 2.34.
Let $X$ be a super $X$-set parameter and let $G$ be a graph of order $n$.
  1. $(1)$
    Then $X(G)\leq\underline{x_{0}}(G)\leq x_{0}(G)$.
  2. $(2)$
    If $G$ has only one minimal $X$-set, then $X(G)=\underline{x_{0}}(G)=x_{0}(G)$.
  3. $(3)$
    If $G$ has more than one minimal $X$-set, then $\overline{X}(G)+1\leq x_{0}(G)\leq\min\{\overline{X}(G)+X(G),n\}$.
  4. $(4)$
    If $G$ has more than one minimum $X$-set, then $X(G)+1\leq\underline{x_{0}}(G)$.
Proof.
(1): This relationship follows immediately from the definitions.
(2): Building onto one minimal $X$-set does not disconnect the graph, therefore $\mathfrak{X}_{k}(G)$ is always connected for $k\geq X(G)$ when there is only one minimal $X$-set.
(3): Let $\hat{S}\subset V(G)$ be minimal $X$-set with $|\hat{S}|=\overline{X}(G)$. Then $\hat{S}$ is an isolated vertex of $\mathfrak{X}_{\overline{X}(G)}(G)$ (because we can’t add a vertex, and removal results in a set that is not an $X$-set). Thus $\overline{X}(G)+1\leq x_{0}(G)$. It is immediate from the definition of $x_{0}(G)$ that $x_{0}(G)\leq n$. Suppose that $\overline{X}(G)+X(G)<n$ and let $k_{0}=\overline{X}(G)+X(G)$. Let $S\subset V(G)$ be a minimal $X$-set of $G$ and $S^{\prime}\subset V(G)$ be a minimum $X$-set of $G$. To ensure $\mathfrak{X}_{k}(G)$ is connected for all $k\geq k_{0}$, it is sufficient to show that every such pair of vertices $S$ and $S^{\prime}$ is connected in $\mathfrak{X}_{k_{0}}(G)$. Define $S^{\prime\prime}=S\cup S^{\prime}$ and observe that $|S^{\prime\prime}|\leq k_{o}$. Then each of $S$ and $S^{\prime}$ is connected by a path to $S^{\prime\prime}$ by adding one vertex at a time. Thus $x_{0}(G)\leq\overline{X}(G)+X(G)$.
(4): Each minimum $X$-set is an isolated vertex in $\mathfrak{X}_{X(G)}(G)$. ∎
For each of the parameters discussed in Sections 3.16, it is easy to find examples of graphs for which $\overline{X}(G)+1=x_{0}(G)$ and $\underline{x_{0}}(G)=x_{0}(G)$. However, strict inequalities are also possible, and examples are presented for each of the parameters discussed. The next result provides some such examples.
Corollary 2.35.
Let $X$ be a super $X$-set parameter. If $G$ has more than one minimal $X$-set and $X(G)=1$, then $x_{0}(G)=\overline{X}(G)+1$
The next result expands and extends Lemma 4 in [19] to super $X$-set parameters; this lemma established Proposition 2.36(2) for domination.
Proposition 2.36.
Let $X$ be a super $X$-set parameter and let $G$ be a graph of order $n$.
  1. (1)
    If for every pair of minimal $X$-sets $M_{1}$ and $M_{2}$, there is a path between $M_{1}$ and $M_{2}$ in $\mathfrak{X}_{k}(G)$, then $\mathfrak{X}_{k}(G)$ is connected.
  2. (2)
    If $k\geq\overline{X}(G)$ and $\mathfrak{X}_{k}(G)$ is connected, then $x_{0}(G)\leq k$.
  3. (3)
    If $|M_{1}\cup M_{2}|\leq k$ for every pair of minimal $X$-sets $M_{1}$ and $M_{2}$, then $x_{0}(G)\leq k$.
Proof.
Suppose first that for every pair of minimal $X$-sets $M_{1}$ and $M_{2}$, there is a path between $M_{1}$ and $M_{2}$ in $\mathfrak{X}_{k}(G)$. Given two $X$-sets $S_{1},S_{2}\in V(\mathfrak{X}_{k}(G))$, each $S_{i}$ contains a minimal $X$-set $M_{i}$. There are paths in $\mathfrak{X}_{k}(G)$ from $S_{1}$ to $M_{1}$, $M_{1}$ to $M_{2}$, and $M_{2}$ to $S_{2}$, so $\mathfrak{X}_{k}(G)$ is connected.
Now assume $k\geq\overline{X}(G)$, $\mathfrak{X}_{k}(G)$ is connected, and $\ell>k$. Since $k\geq\overline{X}(G)$, $\mathfrak{X}_{k}(G)$ contains every minimal $X$-set and there is a path between every pair of minimal $X$-sets in $\mathfrak{X}_{k}(G)$, which is a subgraph of $\mathfrak{X}_{\ell}(G)$. Thus $\mathfrak{X}_{\ell}(G)$ is connected by (1).
If $|M_{1}\cup M_{2}|\leq k$ for every pair of minimal $X$-sets $M_{1}$ and $M_{2}$, then there is a path through $M_{1}\cup M_{2}$ in $\mathfrak{X}_{k}(G)$ for every pair of minimal $X$-sets $M_{1}$ and $M_{2}$ and $x_{0}(G)\leq k$. ∎
It is sometimes useful to exhibit examples where inequalities are not equalities. The next remark (based on Remark 2.10 in [6]) presents conditions under which a graph $G$ satisfies $x_{0}(G)>\overline{X}(G)+1$.
Remark 2.37.
Let $X$ be a super $X$-set parameter and let $G$ be a graph that has at least two minimal $X$-sets. Suppose $G$ has a minimal $X$-set $S$ such that $|S|=\overline{X}(G)$ and for every $u$ and $v$ such that $u\in S$ and $v\in V(G)\setminus S$, the set $B=(S\setminus\{u\})\cup\{v\}$ is not an $X$-set of $G$. Since there is another minimal $X$-set that is not $S$ and it must be in another component of $\mathfrak{X}_{\overline{X}(G)+1}(G)$, $x_{0}(G)>\overline{X}(G)+1$.
The next result is a generalization of Proposition 2.11 in [6]. Just as Proposition 2.11 was used to bound $pd_{0}(H)$ for an example $H$ such that $pd_{0}(G)>\overline{\gamma}_{p}(H)+1$, Proposition 2.38 is used in the proof of Proposition 5.9 to show that $z^{-}_{0}(H(r))>\overline{\operatorname{Z}}_{-}(H(r))+1$.
Proposition 2.38.
Let $X$ be a super $X$-set parameter. Let $G$ be a graph such that there is a partition of the minimal $X$-sets into two sets $\{S_{1},\dots,S_{k}\}$ and $\{T_{1},\dots,T_{\ell}\}$. Let $s=\max_{i=1}^{k}|S_{i}|$, $t=\max_{j=1}^{\ell}|T_{j}|$, and $p=\max_{i=1,j=1}^{k,\ell}|S_{i}\cap T_{j}|$. Then
  1. 1.
    $x_{0}(G)\geq s+t-p$.
  2. 2.
    If $|M_{1}\cup M_{2}|\leq s+t-p$ for any minimal $X$-sets $M_{1}$ and $M_{2}$, then $x_{0}(G)=s+t-p$.
Proof.
Let $S\in\{S_{1},\dots,S_{k}\}$ with $|S|=s$, and $T\in\{T_{1},\dots,T_{\ell}\}$ with $|T|=t$. Since $|S\cap T|\leq p$, $\operatorname{dist}_{\mathfrak{X}(G)}(S,T)\geq s+t-p$. Thus there is no path between $S$ and $T$ in $\mathfrak{X}_{k}(G)$ for $k$ with $\max\{s,t\}\leq k\leq s+t-p-1$.
Now assume $|M_{1}\cup M_{2}|\leq s+t-p$ for any minimal $X$-sets $M_{1}$ and $M_{2}$. Then $x_{0}(G)\leq s+t-p$ by Proposition 2.36. ∎
Twin vertices were used in [7] to expand one graph $G$ with $\underline{z_{0}}(G)<z_{0}(G)$ to a family of graphs with that property. This twin method can be extended to super $X$-set parameters that satisfy additional conditions related to twins. Vertices $u$ and $w$ in a graph $G$ are called independent twins if $N_{G}(u)=N_{G}(w)$. A set of independent twins is a set $\{u_{1},\dots,u_{r}\}\subseteq V(G)$ such that $u_{i}$ and $u_{j}$ are independent twins for all $1\leq i<j\leq r$. For a graph $G$ with a set of independent twins $T$, define $G_{u}=G-u$ for $u\in T$. The next remark and lemma require a property that is parameter specific (this property is established below for skew forcing in Proposition 5.10).
Definition 2.39.
A super $X$-set parameter $X$ has the twins property if it satisfies the following conditions for every graph $G$ and set of independent twins $T$ with $|T|\geq 3$:
  1. 1.
    Any $X$-set must contain at least $|T|-1$ elements of $T$.
  2. 2.
    If $S_{u}$ is an $X$-set of $G_{u}$, then $S=S_{u}\cup\{u\}$ is an $X$-set of $G$.
  3. 3.
    If $S$ is an $X$-set of $G$ and $u\in S$, then $S_{u}=S\setminus\{u\}$ is an $X$-set of $G_{u}$.
Remark 2.40.
Suppose $X$ is a super $X$-set parameter that satisfies the twins property and $G$ is a graph with a set $T$ of at least three independent twins. Then there is a bijection between $X$-sets of $G_{u}$ and $X$-sets of $G$ that contain $u$, and an $X$-set $S_{u}$ of $G_{u}$ is minimal if and only if $S=S_{u}\cup\{u\}$ is a minimal $X$-set of $G$. Furthermore, $X(G-u)=X(G)-1$ and $\overline{X}(G-u)=\overline{X}(G)-1$ for $u\in T$.
The proof of the next result is very similar to that of Lemma 3.8 in [7] and is omitted.
Lemma 2.41.
Let $X$ be a super $X$-set parameter such that has the twins property. Let $G$ be a graph that has a set of twins $T$ with $|T|\geq 3$. If ${\mathfrak{X}}_{k}(G_{u})$ is connected for some $u$, then ${\mathfrak{X}}_{k+1}(G)$ is connected.
Proof.
For $u\in T$, the graphs $G_{u}$ are isomorphic and hence the ${\mathfrak{X}}_{k}(G_{u})$ are isomorphic as well. Thus if ${\mathfrak{X}}_{k}(G_{u})$ is connected for some $u\in T$, then ${\mathfrak{X}}_{k}(G_{w})$ is connected for all $w\in T$. For $S\subseteq V(G)$ with $u\in S$, let $S_{u}=S\setminus\{u\}$.
Since $r\geq 3$ and $X$ has the twins property, any $X$-set $S$ of $G$ must contain at least two vertices in $T$, say $u$ and $w$. A set $S\subseteq V(G)$ that contains $u,w$ is an $X$-set of $G$ if and only if $S\setminus\{u\}$ is an-$X$-set of $G_{u}$ and $S\setminus\{w\}$ is an-$X$-set of $G_{w}$.
Let $S,S^{\prime}$ be two $X$-sets of $G$ of size $k+1$ or less. Since $r\geq 3$ and each can omit at most one vertex in $T$, their intersection must contain at least one $u\in T$. Then $S\setminus\{u\},S^{\prime}\setminus\{u\}$ are $X$-sets for $G_{u}$. Since by hypothesis ${\mathfrak{X}}_{k}(G_{u})$ is connected, that means that there is a path between $S\setminus\{u\}$ and $S^{\prime}\setminus\{u\}$ in ${\mathfrak{X}}_{k}(G_{u})$ and hence a path between $S,S^{\prime}$ in ${\mathfrak{X}}_{k+1}(G)$. ∎
The proof of the next result uses ideas from Proposition 3.9 in [7], but that result is specific to one graph family (in addition to being stated for zero forcing), so we provide the brief proof here.
Proposition 2.42.
Suppose $X$ is a super $X$-set parameter that has the twins property and $G$ is a graph such that $G$ has a set $T$ of $t\geq 2$ twin vertices, $X(G)<\overline{X}(G)$, and $\mathfrak{X}(G)_{X(G)+1}$ is connected. Define $G(t)=G$ and for $r\geq t$, construct $G(r+1)$ from $G(r)$ by adding one additional independent twin of a vertex in $T$. Then $X(G(r))=X(G)+(r-t)$, $\overline{X}(G(r))=\overline{X}(G)+(r-t)$, $\underline{x_{0}}(G(r))=X(G(r))+1$, and $x_{0}(G(r))>\underline{x_{0}}(G(r))$.
Proof.
The proof is by induction. The base case for induction is $r=t$ and is assumed. For $r>t$, apply Remark 2.40 to show that $G(r)$ has minimal $X$-sets of sizes $X(G(r))=X(G)+(r-t)$ and $\overline{X}(G(r))=\overline{X}(G)+(r-t)$; let $k=X(G)+(r-t)$ and $\ell=\overline{X}(G)+(r-t)$. Apply Lemma 2.41 to show ${\mathfrak{X}}_{k+1}(G(r))$ is connected. Since $\overline{X}(G(r))>X(G(r))$, $G(r)$ has at least two minimal $X$-sets, implying ${\mathfrak{X}}_{k}(G(r))$ and ${\mathfrak{X}}_{\ell}(G(r))$ are disconnected by Proposition 2.34. Thus $\underline{x_{0}}(G(r))=k+1$ and $x_{0}(G(r))>\underline{x_{0}}(G(r))$. ∎
Token jumping (TJ) reconfiguration (also called token exchange), which involves exchanging a vertex between sets of vertices of the same size, has been studied for some cohesive vertex set parameters, particularly for minimum $X$-sets (see, e.g., [17] for standard zero forcing and [6] for power domination). Kamiński, Medvedev, and Milanič study token jumping reconfiguration of independent sets and establish the equivalence of the connectedness of the token jumping reconfiguration graph of independent sets of $G$ of size $k$ and the connectedness of the independence $(k-1)$-TAR graph of $G$ in [30]. Here we show this equivalence extends to super $X$-set parameters (and later present the more obvious extension to sub $Y$-set parameters in Proposition 8.7).
Definition 2.43.
Let $W$ be a cohesive vertex set parameter and let $G$ be a graph. The $k$-token jumping or $k$-TJ reconfiguration graph of $G$ for $W$ takes as vertices the $W$-sets of size $k$, with an edge between two $W$-sets $R_{1}$ and $R_{2}$ if and only if $R_{2}$ can be obtained from $R_{1}$ by exchanging exactly one vertex.
Proposition 2.44.
Let $X$ be a super $X$-set parameter, let $G$ be a graph, and let $S_{1}$ and $S_{2}$ be $X$-sets of $G$ with $|S_{i}|=k,i=1,2$. Then there is a path between $S_{1}$ and $S_{2}$ in the $k$-TJ reconfiguration graph of $G$ if and only if there is a path between $S_{1}$ and $S_{2}$ in $\mathfrak{X}_{k+1}(G)$.
Proof.
If $(S_{1}=R_{1},R_{2},\dots,R_{r-1},R_{r}=S_{2})$ is a path in $k$-TJ reconfiguration graph of $G$, then $(S_{1}=R_{1},R_{1}\cup R_{2},R_{2},R_{2}\cup R_{3},R_{3},\dots,R_{r-1},R_{r-1}\cup R_{r},R_{r}=S_{2})$ a path between $S_{1}$ and $S_{2}$ in $\mathfrak{X}_{k+1}(G)$.
We show that any path in between $S_{1}$ and $S_{2}$ can be replaced by a path that uses only $X$-sets of sizes $k$ and $k+1$. Such a path starts with a $k$-set and must alternate additions and deletions. Each addition-deletion pair can then be replaced by an edge in the $k$-TJ graph of $G$.
So suppose $(S_{1}=R_{1},R_{2},\dots,R_{r-1},R_{r}=S_{2})$ is a path between $S_{1}$ and $S_{2}$ in $\mathfrak{X}_{k+1}(G)$. If $|R_{i}|\geq k$ for $i=1,\dots,r$ then there is nothing to show. So assume that is not the case and let $j$ be an index such that $|R_{j}|$ is minimized. Necessarily $|R_{j}|=|R_{j-1}|-1$, $R_{j}$ is obtained from $R_{j-1}$ by deleting some vertex $w_{j-1}$, $|R_{j+1}|=|R_{j}|+1$, and $R_{j+1}$ is obtained from $R_{j}$ by adding some vertex $u_{j}$. Then replace $R_{j}$ by $R^{\prime}_{j}=R_{j-1}\cup\{u_{j}\}$ so $R^{\prime}_{j}$ is obtained from $R_{j-1}$ by adding $u_{j}$ and $R_{j+1}$ is obtained from $R_{j}$ by deleting $w_{j-1}$. Since $X$ is a super $X$-set parameter and $R_{j-1}$ is an $X$-set, necessarily $R^{\prime}_{j}$ is an $X$-set. This process can be repeated as needed to obtain a path in which all $X$-sets have size $k$ or $k+1$. ∎
2.6 $X$-TAR graph uniqueness
Let $X$ be a robust $X$-set parameter. If $X(K_{1})=1$, for a graph $H$ with no isolated vertices, we say its TAR graph is unique if $\mathfrak{X}(G)\cong\mathfrak{X}(H)$ implies $G\cong H$ for any graph $G$ with no isolated vertices. If $X(K_{1})=0$, for a graph $H$, we say its TAR graph is unique if $\mathfrak{X}(G)\cong\mathfrak{X}(H)$ implies $G\cong H$ for any graph $G$. The concept of unique TAR graphs was introduced in [6] for power domination. Certainly whether a graph has a unique TAR graph is parameter specific, and is discussed for various parameters in Sections 36. For power domination and standard zero forcing, examples of graphs with unique TAR graphs were presented in [6] and [7]. The fact that the order of $G$, $X(G)$, and $\overline{X}(G)$ can all be determined from $\mathfrak{X}(G)$ was used in [6], and the fact that $\mathfrak{X}(G)\cong\mathfrak{X}(H)$ implies we can relabel so $G$ and $H$ have the same $X$-sets was essential to the uniqueness results in [7]. This ability to assume that $G$ and $H$ have the same $X$-sets is true for robust $X$-set parameters, and we use this to present examples of unique TAR graphs for domination, PSD forcing, skew forcing, and vertex covers in Sections 3.1, 4.1, 5.1, and 6.1. We also present data for the frequency of unique TAR graphs among small graphs in those sections. We point out that vertex cover TAR graphs behave quite differently from the TAR graphs of the other parameters discussed here: it is shown in Proposition 6.6 that every vertex cover TAR graph is unique. Which specific graphs have unique TAR graphs naturally depends on the parameter, but there are some graph families that have unique TAR graphs for many of the parameters discussed, such as complete graphs and complete bipartite graphs. Even when the same graph family has unique TAR graphs for multiple parameters, the proofs are usually parameter-specific (although some ideas recur).
2.7 Hamilton paths and cycles and cut-vertices in TAR graphs
The study of whether a TAR graph has a Hamilton path was initiated in [1] for domination. As shown there, no domination TAR graph can have a Hamilton cycle (because any cycle in a bipartite graph is even and there are an odd number of dominating sets), but for some of the other parameters, such as skew forcing, it is possible to have a Hamilton cycle.
Remark 2.45.
Let $X$ be a super $X$-set parameter. It is well-known that a hypercube of dimension at least two has a Hamilton cycle. Thus $\mathfrak{X}(G)$ has a Hamilton cycle whenever $\mathfrak{X}(G)\cong Q_{n}$, which happens if the empty set is an $X$-set. The graph $Q_{n}-v$ for $v\in V(Q_{n})$ is realized as a TAR graph whenever any one vertex of $G$ is an $X$-set. Since deleting a vertex from a cycle leaves a path and $Q_{n}-v$ has an odd number of vertices, $Q_{n}-v$ has an Hamilton path but not a Hamilton cycle.
Lemma 6.7 in [1] shows that $V(G)$ is the only possible cut-vertex of the domination TAR graph of $G$ for any graph $G$. We extend this from domination to super $X$-set parameters.
Proposition 2.46.
Let $X$ be a super $X$-set parameter and let $G$ be a graph of order $n$. If $S$ is a cut-vertex of $\mathfrak{X}(G)$, then $S=V(G)$.
Proof.
We prove the contrapositive. Assume that $S\neq V(G)$. If $S$ has exactly one neighbor, then $S$ is not a cut-vertex. So, suppose that $S$ has distinct neighbors $A,B\subseteq V(G)$ in $\mathfrak{X}(G)$. It suffices to show that there is a path in $\mathfrak{X}(G)-S$ between $A$ and $B$. Since $S\neq V(G)$, there is some vertex $w$ of $G$ such that $w\not\in S$. Define $S^{\prime}=S\cup\{w\}$, $A^{\prime}=A\cup\{w\}$ and $B^{\prime}=B\cup\{w\}$. Note that $S^{\prime}\neq S$. If $w\not\in A$ then $A\neq A^{\prime},S^{\prime}$ and $A^{\prime}\neq S^{\prime}$, and similarly for $B$. Thus if $w\not\in A$ and $w\not\in B$, then $(A,A^{\prime},S^{\prime},B^{\prime},B)$ is a path from $A$ to $B$ in $\mathfrak{X}(G)-S$. If $w\in A$ and $w\not\in B$, then $A=A^{\prime}=S^{\prime}$ and $(A,B^{\prime},B)$ is a path from $A$ to $B$ in $\mathfrak{X}(G)-S$, and similarly when $w\not\in A$ and $w\in B$. It is not possible to have $w\in A$ and $w\in B$. ∎
Observe that if $X$ is a connected $X$-set parameter, $G$ is a graph of order $n\geq 2$, and $\overline{X}(G)=n-1$, then $V(G)$ is a cut-vertex of $\mathfrak{X}(G)$ because any minimal $X$-set with $n-1$ vertices is a connected component in $\mathfrak{X}(G)-V(G)$ (and $G$ has more than one minimal $X$-set).
3 TAR reconfiguration results for domination, power domination, and standard zero forcing
TAR graphs have been studied previously for domination, power domination, and zero forcing. In this section we review these results and provide some additional material.
3.1 Domination TAR graphs
As noted in [7], domination is an original $X$-set parameter, so Theorem 2.23 applies. We summarize known results and provide some additional material related to domination TAR graphs, specifically uniqueness, Hamiltonicity, and connectedness results.
The domination TAR graph of a base graph $G$ is denoted by $\mathfrak{D}(G)$ and the upper domination number of $G$ is denoted here by $\overline{\gamma}(G)$ (note that in the literature this parameter is usually dented by $\Gamma(G)$).
Example 3.1.
Since $K_{n}$ is the only graph of order $n$ for which any one vertex dominates, $K_{n}$ is the only graph such that $\mathfrak{D}(K_{n})\cong Q_{n}-v$ for $v\in V(Q_{n})$ and $\mathfrak{D}(K_{n})$ is unique. As was noted in [1], $\mathfrak{D}(K_{n})$ has a Hamilton path but not a Hamilton (see also Remark 2.45).
Example 3.2.
Let $2\leq p\leq q$ and let $A$ and $B$ be the partite sets of $K_{p,q}$. Then the minimal dominating sets of $K_{p,q}$ are $A$, $B$ and $\{a_{i},b_{j}\}$ where $a_{i}\in A$ and $b_{j}\in B$. Thus $\gamma(K_{p,q})=2$ and $\overline{\gamma}(K_{p,q})=q$.
Complete biparite graphs are discussed further in Example 3.11 as well as in the next result.
Proposition 3.3.
For $1\leq p\leq q$, $\mathfrak{D}(K_{p,q})$ is unique unless $q=p\geq 3$. For $p=q\geq 3$, if $\mathfrak{D}(G)\cong\mathfrak{D}(K_{p,p})$, then $G\cong\mathfrak{D}(K_{p,p})$ or $G\cong\mathfrak{D}(K_{p}\Box K_{2})$.
Proof.
Suppose $G$ has no isolated vertices and $\mathfrak{D}(G)\cong\mathfrak{D}(K_{p,q})$. This implies $G$ has the same dominating sets as $K_{p,q}$ (possibly after relabeling). If $p=1$, then $\gamma(G)=1$ and $\overline{\gamma}(G)=q=|V(G)|-1$, which implies $G\cong K_{1,q}$. Thus $\mathfrak{D}(K_{1,q})$ is unique, and we assume $q\geq p\geq 2$.
We begin by showing that this implies $G[B]\cong qK_{1}$ or $G[B]\cong K_{q}$. Suppose that $G[B]\not\cong K_{q}$, i.e., there exist $k,\ell$ such that $b_{k}$ and $b_{\ell}$ are not adjacent. Since $\{a_{i},b_{k}\}$ is a dominating set for $i=1,\dots,p$, $A\subseteq N_{G}[b_{\ell}]$ (and similarly $A\subseteq N_{G}[b_{k}]$). Suppose there is an edge in $G[B]$, i.e., there exist $i,j$ such that $b_{i}$ and $b_{j}$ are adjacent. Necessarily $\{i,j\}\neq\{k,\ell\}$, so without loss of generality $i\neq k,\ell$. Then $B\setminus\{b_{i}\}$ is a dominating set, contradicting the minimal dominating sets of $G$. Thus $G[B]\cong qK_{1}$ or $G[B]\cong K_{q}$. Similarly, $G[A]\cong pK_{1}$ or $G[A]\cong K_{p}$.
Now suppose $G[B]\cong qK_{1}$; we show this implies $G\cong K_{p,q}$. By the argument above, $G[B]\cong qK_{1}$ implies every vertex of $B$ is adjacent to every vertex of $A$. If $a_{i}$ and $a_{j}$ were adjacent, then $A\setminus\{a_{i}\}$ would be a dominating set, which is a contradiction. Thus $G[A]\cong pK_{1}$ and $G\cong K_{p,q}$.
The previous argument applies to $G[A]\cong pK_{1}$ also, so the only remaining case is $G[B]\cong K_{q}$ and $G[A]\cong K_{p}$, and we assume this. We show that for every $k=1,\dots,q$, there must exist $\ell\in\{1,\dots,p\}$ such that $N_{G}[a_{\ell}]\cap B=\{b_{k}\}$. Once this is established, necessarily $q=p$ and $G\cong K_{p}\Box K_{2}$. So suppose not, i.e., there is some $b_{k}$ such that $N_{G}[a_{i}]\cap B\neq\{b_{k}\}$ for every $i=1,\dots,p$. Then $B\setminus\{b_{k}\}$ is a dominating set because $b_{k}$ is dominated by any other vertex of $B$ and $A$ is dominated by $B$. Since this is a contradiction, $G\cong K_{p}\Box K_{2}$.
Let $G\cong K_{p}\Box K_{2}$ with $p\geq 2$, and denote the vertices of one copy of $K_{p}$ by $A$ and the other by $B$, Then minimal dominating sets of $G$ are $A$, $B$ and $\{a_{i},b_{j}\}$ where $a_{i}\in A$ and $b_{j}\in B$. Thus $\mathfrak{D}(G)\cong\mathfrak{D}(K_{p,q})$. Note that for $p\geq 3$, $K_{p}\Box K_{2}\not\cong K_{p,p}$. Since $K_{2}\Box K_{2}\cong K_{2,2}$, $\mathfrak{D}(K_{2,2})$ is unique. ∎
Table 3.1 shows the number of (nonismorphic) graphs without isolated vertices of order at most eight that have unique domination TAR graphs (this data was computed in [11]).
# vertices in $G$ 2 3 4 5 6 7 8
# graphs with unique $\mathfrak{D}(G)$ 1 2 5 14 55 348 4275
# graphs with no isolated vertices 1 2 7 23 122 888 11302
ratio (# unique/# no isolated) 1 1 0.7143 0.6087 0.4508 0.3919 0.3783
Table 3.1: Number of graphs having a unique domination TAR graph for graphs of small order
The question of whether $\mathfrak{D}(G)$ has a Hamilton path is resolved for complete graphs and complete bipartite graphs in [1], and for trees and cycles in [2].
Theorem 3.4.
[1] For $1\leq p\leq q$, $\mathfrak{D}(K_{p,q})$ has a Hamilton path if and only if at least one of $p$ or $q$ is odd.
Theorem 3.5.
[2] For any tree $T$, $\mathfrak{D}(T)$ has a Hamilton path.
Theorem 3.6.
[2] For $n\geq 3$, $\mathfrak{D}(C_{n})$ has a Hamilton path if and only if n is not a multiple of $4$.
Haas and Seyffarth ([19]) were the first to investigate the connectedness of $\mathfrak{D}_{k}(G)$. The least $k_{0}$ such that $\mathfrak{D}_{k}(G)$ is connected for all $k\geq k_{0}$ is denoted by $d_{0}(G)$, and the least $k$ such that $\mathfrak{D}_{k}(G)$ is connected is denoted by $\underline{d_{0}}(G)$. They established the bounds for $d_{0}$ found in Proposition 2.34. It was also shown there that if $G$ is a graph of order $n$ with at least two disjoint edges, then $\mathfrak{D}_{n-1}(G)$ is always connected. They found graph families for which $d_{0}$ is equal to the lower bound, such as the one listed next.
Theorem 3.7.
[19] For any nontrivial bipartite or nontrivial chordal graph, $\mathfrak{D}_{\overline{\gamma}(G)+1}(G)$ is connected.
Conditions that guarantee a graph will satisfy $d_{0}(G)=\overline{\gamma}(G)+1$ were presented in [20]. The first examples of a graph family for which $\mathfrak{D}_{\overline{\gamma}(G)+1}$ is not connected is due to Suzuki, Mouawad and Nishimura in [37]. They constructed a family of graphs $G_{(b,d)}$ from $d+1$ cliques of size $b$ with additional edges such that $d_{0}(G_{(d,b)})=\overline{\gamma}(G_{(d,b)})+2$. The next example shows a planar graph in this family.
111122223333444455556666777788889999
Figure 3.1: The graph $G=K_{3}\Box P_{3}$, which has $\gamma(G)=\overline{\gamma}(G)=3$ and $d_{0}(G)=5>\overline{\gamma}(G)+1$
Example 3.8.
The graph $G_{(2,3)}$ in [37] is isomorphic to $G=K_{3}\Box P_{3}$. With the vertices labelled as shown in Figure 3.1, the minimal dominating sets are {4, 5, 6}, {1, 2, 9}, {2, 6, 9}, {3, 5, 9}, {2, 4, 8}, {3, 4, 8}, {2, 5, 9}, {2, 6, 7}, {3, 5, 8}, {1, 6, 8}, {1, 5, 7}, {1, 8, 9}, {1, 5, 9}, {2, 5, 8}, {2, 5, 7}, {1, 4, 9}, {1, 6, 9}, {2, 7, 9}, {3, 6, 8}, {1, 5, 8}, {3, 6, 9}, {1, 4, 7}, {1, 6, 7}, {2, 3, 7}, {2, 6, 8}, {3, 6, 7}, {1, 3, 8}, {3, 4, 9}, {3, 7, 8}, {1, 4, 8}, {2, 4, 9}, {3, 4, 7}, {2, 4, 7}, {3, 5, 7} (this is documented in [11]). Observe that a minimal dominating set is a set of three vertices that is the vertices of middle $K_{3}$ (the set of vertices $\{4,5,6\}$) or has exactly one element in each $K_{3}$. Thus $\gamma(G)=\overline{\gamma}(G)=3$ and $d_{0}(G)=5>\overline{\gamma}(G)+1$.
For any positive integer $k$, a $k$-tree is a graph that can be constructed from $K_{k+1}$ by repeatedly (possibly zero times) adding a new vertex and joining it to an existing $k$-clique. Thus, a $1$-tree is a tree, and a tree of order two or more is a 1-tree. For a graph $G$, the minimum $k$ such that $G$ is a subgraph of some $k$-tree is called the tree-width of $G$. A graph is $b$-partite if the vertex set can be partitioned into sets $V_{1},\dots,V_{b}$ such that for every edge the end points are in distinct partite sets.
Theorem 3.9.
[37] For every integer $b\geq 3$ there exists an infinite family of graphs of tree-width $12b-$ 1 such that for each $G$ in the family $\mathfrak{D}_{\overline{\gamma}(G)+1}(G)$ is not connected. There also exists an infinite family of $b$-partitie graphs such that $\mathfrak{D}_{\overline{\gamma}(G)+1}$ is not connected for any $G$ in the family.
In [33], Mynhardt, Teshima and Roux constructed graphs with arbitrary domination number and arbitrary upper domination number at least three and greater than domination number that realize the upper bound for $d_{0}(G)$, as stated in the next theorem. This construction was built on another construction in [33] that realizes $d_{0}(G)=\gamma(G)+\overline{\gamma}(G)-1$.
Theorem 3.10.
[33] For each integer $k\geq 3$ and each integer $r$ such that $11\leq r\leq k-$ 1, there is a graph $Q_{k,r}$ such that $\gamma(Q_{k,r})=r$, $\overline{\gamma}(Q_{k,r})=k$ and $d_{0}(Q_{k,r})=k+r=\gamma(Q_{k,r})+\overline{\gamma}(Q_{k,r})$.
The graph $K_{n}$ has $d_{0}(K_{n})=\underline{d_{0}}(K_{n})$. The next graph is an example where $\underline{d_{0}}(G)<d_{0}(G)$.
Example 3.11.
Let $4\leq p\leq q$ and let $A$ and $B$ be the partite sets of $K_{p,q}$. As noted in Example 3.2, the minimal dominating sets of $K_{p,q}$ are $A$, $B$ and $\{a_{i},b_{j}\}$. Thus $\underline{d_{0}}(K_{p,q})=3$ and $d_{0}(K_{p,q})=q+1\geq 5$.
For more information on domination TAR reconfiguration see the survey by Mynhardt and Nasserasr [32] and the references therein, as well as to [19, 20, 33, 37].
3.2 Power domination TAR graphs
Token addition and removal reconfiguration for power domination and token jumping reconfiguration of minimum power dominating sets were studied in [6], where the first definition of an original $X$-set property was given and results were established for TAR graphs of original $X$-set parameters. As noted in [7], power domination is an original $X$-set parameter, so Theorem 2.23 applies. Here we summarize parameter-specific results for power domination TAR graphs from [6] and [7], and present some additional information. The power domination TAR graph of a base graph $G$ is denoted by $\mathfrak{P}(G)$ and the upper power domination number of $G$ is denoted by $\overline{\gamma}_{p}(G)$.
Example 3.12.
It is shown in [6, Section 3.3] that $\mathfrak{P}(K_{1,q})$ is unique for $q\geq 3$ but $\mathfrak{P}(K_{2,q})$ is not for $q\geq 2$.
It is also proved in [6] that $\mathfrak{P}(K_{3,3})$ is unique, and established computationally that $\mathfrak{P}(K_{3,4})$, $\mathfrak{P}(K_{3,5})$, $\mathfrak{P}(K_{4,4})$ and $\mathfrak{P}(K_{4,5})$ are unique. It is conjectured there that $\mathfrak{P}(K_{p,q})$ is unique for $q\geq p\geq 3$. Table 3.2 shows the number of (nonisomorphic) graphs without isolated vertices of order at most eight that have unique power domination TAR graphs (this data was computed in [11]).
# vertices in $G$ 2 3 4 5 6 7 8
# graphs with unique $\mathfrak{P}(G)$ 1 0 3 4 13 25 79
# graphs with no isolated vertices 1 2 7 23 122 888 11302
ratio (# unique/# no isolated) 1 1 0.4286 0.1739 0.1066 0.0282 0.00699
Table 3.2: Number of graphs having a unique power domination TAR graph for graphs of small order
Next we examine the Hamiltonicity or lack thereof for power domination TAR graphs, starting with examples from [6].
Example 3.13.
If $G$ is a graph such that any one of its vertices is a power dominating set, then $\mathfrak{P}(G)$ is isomorphic to an $n$-dimensional hypercube with one vertex (the empty set) deleted. Examples of such base graphs include $C_{n},P_{n},K_{n},$ and the wheel $W_{n}=C_{n-1}\vee K_{1}$ [6]. Thus the power domination TAR graphs $\mathfrak{P}(G)$ for $G=C_{n},P_{n},K_{n},W_{n}$ each have a Hamilton path but not a Hamilton cycle.
Example 3.14.
The graph $G(r,1)$ is constructed by adding one leaf vertex $\ell$ to $K_{r}$ and it is shown in [7, Proposition 2.16] that the vertex $\ell$ is $\gamma_{p}$-irrelevant. Thus $\mathfrak{P}(G(r,1))\cong(Q_{r}-v)\Box K_{2}$ where $v\in V(Q_{r})$, which has a Hamilton cycle because $Q_{r}-v$ has a Hamilton path.
Finally, we present connectedness examples. The least $k_{0}$ such that $\mathfrak{P}_{k}(G)$ is connected for all $k\geq k_{0}$ is denoted by $pd_{0}(G)$ and the least $k$ such that $\mathfrak{P}_{k}(G)$ is connected is denoted by $\underline{pd_{0}}(G)$. There are many examples of graphs $G$ where $pd_{0}(G)=\overline{\gamma}_{p}(G)+1$ and $pd_{0}(G)=\underline{pd_{0}}(G)$, including any graph $G$ of order $n$ that has $\mathfrak{P}(G)\cong Q_{n}-v$ for $v\in V(Q_{n})$. An example from [6] that shows $pd_{0}(G)>\overline{\gamma}_{p}(G)+1$ is possible and a new example that shows $pd_{0}(G)>\underline{pd_{0}}(G)$ is possible. For any integer $n\geq 3$, let $G_{n}=(V_{n},E_{n})$ be the graph defined as follows: $V_{n}=T_{n}\cup\big{(}\cup_{i=1}^{n}S_{n,i}\big{)}$ where $T_{n}=\{u_{1},\ldots,u_{n-1}\}$ and $S_{n,i}=\{v^{i}_{1},\ldots,v^{i}_{n}\}$, for every integer $i$, $1\leq i\leq n$. To define $E_{n}$: For integers $i=1,\ldots,n$ and $j=1,\ldots,n-1$, $N_{G_{n}}(v^{i}_{j})=\{u_{j}\}\cup\{v^{i}_{p}:1\leq p\leq n,p\not=j\}$; $N_{G_{n}}(v^{i}_{n})=\{v^{i}_{1},\ldots,v^{i}_{n-1}\}$; $N_{G_{n}}(u_{j})=\{v^{r}_{j}:1\leq r\leq n\}$ [6, Definition 3.11].
Theorem 3.15.
[6, Theorem 3.12] For $n\geq 3$, the graph $G_{n}$ has $\overline{\gamma}_{p}(G_{n})=\gamma_{p}(G_{n})=n-1$ and $\underline{pd_{0}}(G_{n})=pd_{0}(G_{n})=2n-2>\overline{\gamma}_{p}(G_{n})+1=\gamma_{p}(G_{n})+1$.
Next we present an example of a graph $G$ having $pd_{0}(G)>\underline{pd_{0}}(G)$. For $H$ a connected graph of order at least two and $q\geq 3$, define the graph $K^{2,q}(H)$ to be the graph obtained from $H$ deleting each edge of $H$ and adding $q$ vertices each adjacent exactly to the endpoints of the edge that was deleted. Keep the vertex labels of $H$ in $K^{2,q}(H)$, so $V(H)\subseteq V(K^{2,q}(H))$ (this extends the definition of a family of graphs introduced in [6] in the study of token exchange reconfiguration). The next result was established for $q=3$ in Lemma 5.6 of [6] and the proof remains valid for $q\geq 3$.
Lemma 3.16.
Let $H$ be a connected graph of order at least two and let $q\geq 3$. If $S$ is a minimum power dominating set of $K^{2,q}(H)$, then $S\subseteq V(H)$. A set $S\subseteq V(H)$ is a minimum vertex cover of $H$ if and only if $S$ is a minimum power dominating set of $K^{2,q}(H)$.
Proof.
Denote the vertices of $H$ by $v_{1},\dots,v_{n}$ and note that $\{v_{1},\dots,v_{n}\}\subseteq V(K^{2,q}(H))$. Suppose first that $S$ is a power dominating set of $K^{2,q}(H)$. For every $i,j\in\{1,\dots,n\}$ with $v_{i}v_{j}\in E(H)$, one of $v_{i}$ or $v_{j}$ must be in $S$ or at least $q-1$ of the $q$ vertices in $N_{K^{2,q}(H)}(v_{i})\cap N_{K^{2,q}(H)}(v_{j})$ must be in $S$. Observe that $q-1\geq 2$. Thus in order for $S$ to be a minimum power dominating set, $S$ must contain one of $v_{i}$ or $v_{j}$ for every $v_{i}v_{j}\in E(H)$ (and no vertices in $V(K^{2,q}(H))\setminus V(H)$). Thus $S$ is a vertex cover of $H$.
Now suppose $S$ is a minimum vertex cover of $H$ and let $v_{i}v_{j}\in E(H)$ with $v_{i}\in S$. Then the set of $q$ vertices in $N_{K^{2,q}(H)}(v_{i})\cap N_{K^{2,q}(H)}(v_{j})$ is dominated and any one of these vertices can force $v_{j}$ if $v_{j}\not\in S$. Thus $S$ is a power dominating set of $K^{2,q}(H)$. Since every minimum power dominating set of $K^{2,q}(H)$ is a subset of $V(H)$, $S$ is a minimum power dominating set of $K^{2,q}(H)$. ∎
Proposition 3.17.
For $r\geq 3$ and $q\geq 4$, $\gamma_{p}(K^{2,q}(K_{r}))=r-1$, $\overline{\gamma}_{p}(K^{2,q}(K_{r}))\geq q\left(\frac{r(r-1)}{2}\right)-(r-1)$, $\underline{pd_{0}}(K^{2,q}(K_{r}))=r$, and $pd_{0}(K^{2,q}(K_{r}))\geq q\left(\frac{r(r-1)}{2}\right)-(r-1)+1>r=\underline{pd_{0}}(K^{2,q}(K_{r}))$.
Proof.
By Lemma 3.16, $\gamma_{p}(K^{2,q}(K_{r}))=r-1$ and any set of $r-1$ vertices of $K_{r}$ is a minimum power dominating set. Since each of these is adjacent to the vertex $V(K_{r})$ in $\mathfrak{P}_{r}(K^{2,q}(K_{r}))$ and the only minimal power dominating sets of at most $r$ vertices are the minimum power dominating sets, $\mathfrak{P}_{r}(K^{2,q}(K_{r}))$ is connected and $\underline{pd_{0}}(K^{2,q}(K_{r}))=r$.
Next we exhibit a minimal power dominating set containing of $q\left(\frac{r(r-1)}{2}\right)-(r-1)$ vertices, which shows that $\overline{\gamma}_{p}(K^{2,q}(K_{r}))\geq q\left(\frac{r(r-1)}{2}\right)-(r-1)$ and $pd_{0}(K^{2,q}(K_{r}))\geq q\left(\frac{r(r-1)}{2}\right)-(r-1)+1$. It can be verified algebraically that $q\left(\frac{r(r-1)}{2}\right)-(r-1)+1>r$ for $r\geq 3$ and $q\geq 4$.
Denote the vertices of $K_{r}$ by $v_{1},\dots,v_{r}$ and denote the $q$ vertices of $K^{2,q}(K_{r})$ obtained by replacing the edge $v_{i}v_{j}$ by $u^{i,j}_{k}$ for $k=1,\dots,q$. Define $S$ to be the set of the first $q-1$ degree-two vertices associated with each of the $r-1$ edges incident with $v_{1}$ and all $q$ degree-two vertices associated with each of the edges not incident with $v_{1}$, i.e., $S=\{u^{1,j}_{k}:k=1,\dots,q-1,j=2,\dots,r\}\cup\{u^{i,j}_{k}:k=1,\dots,q,j=i+1,\dots,r,i=2,\dots,r\}$. Observe that $|S|=q\left(\frac{r(r-1)}{2}\right)-(r-1)$. Starting with the vertices in $S$ blue, after the first (domination) step all vertex are blue except $v^{1,j}_{q}$ for $j=2,\dots,r$. Then $v_{j}$ can observe $v^{1,j}_{q}$ in the next round. ∎
3.3 Standard zero forcing TAR graphs
TAR reconfiguration was studied for standard zero forcing in [7]. The main isomorphism results were proved for original $X$-sets there. Both those results and results for original $X$-set parameters were applied to standard zero forcing. Additional parameter-specific results for standard zero forcing TAR graphs are also presented in [7]. We briefly summarize the latter results here, and also examine the Hamiltonicity of some examples of standard zero forcing TAR graphs from [7]. The standard zero forcing TAR graph of a base graph $G$ is denoted by $\mathfrak{Z}(G)$ and the upper standard zero forcing number of $G$ is denoted here by $\overline{\operatorname{Z}}(G)$
As noted in [7, Section 3.1], $K_{n}$ is the only graph $G$ of order $n$ that has no isolated vertices and for which $\operatorname{Z}(G)=n-1$, so $\mathfrak{Z}(K_{n})$ is unique. Similarly, $P_{n}$ is the only graph $G$ of order $n$ that has no isolated vertices and for which $\operatorname{Z}(G)=1$, so $\mathfrak{Z}(P_{n})$ is unique. It is also shown there that $\mathfrak{Z}(K_{1,q})$ is unique for $q\geq 2$ and that $\mathfrak{Z}(C_{n})$ is not unique for $n\geq 4$. For the convenience of the reader in comparing across parameters, we reproduce the table of uniqueness data for standard zero forcing TAR graphs of small base graphs.
# vertices in $G$ 2 3 4 5 6 7 8
# graphs with unique $\mathfrak{Z}(G)$ 1 2 4 7 34 303 5318
# graphs with no isolated vertices 1 2 7 23 122 888 11302
ratio (# unique/# no isolated) 1 1 0.5714 0.3043 0.2787 0.3412 0.4705
Table 3.3: [7, Table 1] Number of graphs with unique standard zero forcing TAR graph for small orders
Example 3.18.
It is shown in [7] that $\mathfrak{Z}(K_{n})\cong K_{1,n}$ and $\mathfrak{Z}(K_{1,q})\cong K_{1,q}\Box K_{2}$. Thus $\mathfrak{Z}(K_{n})$ does not have a Hamilton path for $n\geq 3$. Similarly, $\mathfrak{Z}(K_{1,q})$ has a Hamilton cycle if and only if $q\leq 2$ if and only if $\mathfrak{Z}(K_{1,q})$ has a Hamilton path.
Example 3.19.
Observe that if $H$ is a graph of even order with a Hamilton cycle $C$, then $H\Box P_{r}$ also has a Hamilton cycle, by alternating edges of $C$ in the first and last copies of $H$ and moving between these two cycles on copies of $P_{r}$. Note that $\mathfrak{Z}(K_{1,2})$ has a Hamilton cycle and is of even order and $\mathfrak{Z}(K_{2})\cong P_{3}$ (Example 3.18). So $\mathfrak{Z}(K_{1,2}\,\sqcup\,K_{2}\dots\,\sqcup\,K_{2})\cong\mathfrak{Z}(K_{1,2}\Box P_{3}\dots\Box P_{3}$) by Proposition 2.8. Hence there are arbitrarily large order base graphs $K_{1,2}\,\sqcup\,K_{2}\dots\,\sqcup\,K_{2}$ that have standard zero forcing TAR graphs with Hamilton cycles.
The least $k_{0}$ such that $\mathfrak{Z}_{k}(G)$ is connected for all $k\geq k_{0}$ is denoted by $z_{0}(G)$ and the least $k$ such that $\mathfrak{Z}_{k}(G)$ is connected is denoted by $\underline{z_{0}}(G)$. It is common to have $z_{0}(G)=\overline{\operatorname{Z}}(G)+1$ and $\underline{z_{0}}(G)=z_{0}(G)$, as illustrated by $K_{n}$ and in the next example.
Example 3.20.
[7] Number the vertices of the path $P_{n}$ with $n\geq 4$ in path order. A set $S\subseteq V(P_{n})$ is a standard zero forcing set if and only if $S$ contains an endpoint or $S$ contains two consecutive vertices of the path. The set $\{2,3\}$ is a standard zero forcing set, but is not adjacent to any zero forcing set in $\mathfrak{Z}_{2}(P_{n})$. Thus $\overline{\operatorname{Z}}(G)=2$ and $\mathfrak{Z}_{2}(P_{n})$ is not connected. Adding an end vertex makes any set a standard zero forcing set, so $\mathfrak{Z}_{k}(G)$ is connected for $k\geq 3$. Thus $\underline{z_{0}}(P_{n})=3=z_{0}(P_{n})=\overline{\operatorname{Z}}(P_{n})+1$.
Next we present connectedness examples from [7] showing $z_{0}(G)>\overline{\operatorname{Z}}(G)+1$ and $z_{0}(G)>\underline{z_{0}}(G)$ are possible. The graph $H(r)$ is constructed from two copies of $K_{r+2}$ by adding a matching between $r$ pairs of vertices.
Proposition 3.21.
[7, Proposition 3.4] For $r\geq 2$, $\operatorname{Z}(H(r))=\overline{\operatorname{Z}}(H(r))=r+2$ and $\underline{z_{0}}(H(r))=z_{0}(H(r))=2r+2=\overline{\operatorname{Z}}(H(r))+r$.
The next family of graphs provides examples of graphs $G$ such that $z_{0}(G)>\underline{z_{0}}(G)$. For $r\geq 3$, construct $H_{r}$ from $H_{2}$ (shown in Figure 3.2) by adding vertices $u_{3},\dots,u_{r}$ with $N_{H_{r}}(u_{k})=N_{H_{r}}(u_{1})$ for $k=3,\dots,r$ [7]. Note that $|V(H_{r})|=r+6$. The next result is established by use of the twins property (although it is not called that in [7]).
$u_{1}$$u_{2}$
Figure 3.2: A graph $H_{2}$ satisfying $z_{0}(H_{2})>\underline{z_{0}}(H_{2})$
Proposition 3.22.
[7, Proposition 3.9] For $r\geq 2$, $\operatorname{Z}(H_{r})=r+2$, $\overline{\operatorname{Z}}(H_{r})=r+4$, $\underline{z_{0}}(H_{r})=r+3$, and $z_{0}(H_{r})=r+5$.
4 TAR reconfiguration for positive semidefinite (PSD) forcing
It was noted in [7] that PSD forcing is an original $X$-set parameter, so the original $X$-set parameter results in [6, 7] apply to PSD forcing. In particular, Theorem 2.23 applies to PSD forcing. It was noted there that there are no irrelevant vertices for PSD forcing [7, Proposition 2.19]. Here we determine PSD forcing TAR graphs (also called $\operatorname{Z}_{+}$-TAR graphs) for certain graph families, discuss uniqueness, Hamiltonicity, and connectedness of $\operatorname{Z}_{+}$-TAR graphs. Examples showing connectedness bounds need not be equalities are presented. Denote the $\operatorname{Z}_{+}$-TAR graph of a base graph $G$ by $\mathfrak{Z}^{+}(G)$ and denote the upper PSD forcing number of $G$ by $\overline{\operatorname{Z}}_{+}(G)$.
4.1 Examples of PSD TAR graphs, uniqueness, and Hamiltonicity
We present examples of $\operatorname{Z}_{+}$-TAR graphs that illustrate both uniqueness and nonuniqueness and the existence of graphs with Hamilton cycles, Hamilton paths but not cycles, and not having Hamilton paths.
Example 4.1.
Since $\operatorname{Z}_{+}(K_{n})=n-1$ and $K_{n}$ is the only graph of order $n$ with this property, $\mathfrak{Z}^{+}(K_{n})\cong K_{1,n}$ and $K_{n}$ has a unique $\operatorname{Z}_{+}$-TAR graph. Thus $\mathfrak{Z}^{+}(K_{n})$ does not have a Hamilton path for $n\geq 3$.
Example 4.2.
For a tree $T$ of order $n$, any one vertex is a PSD zero forcing set (and the empty set is not a PSD forcing set of any graph). Thus $\mathfrak{Z}^{+}(T)\cong Q_{n}-v$ for $v\in V(Q_{n})$ (cf. Remark 2.18). Thus $\mathfrak{Z}^{+}(T)$ is not unique for $n\geq 4$. Furthermore, $\mathfrak{Z}^{+}(T)$ has a Hamilton path for every tree.
Example 4.3.
For a cycle $C_{n}$, any set of two vertices is a PSD zero forcing set and $\operatorname{Z}_{+}(C_{n})=2$. Thus $\mathfrak{Z}^{+}(C_{n})$ is isomorphic to the graph obtained from $Q_{n}$ by removing the bottom two layers (corresponding to sets of cardinality zero and one). That is, $\mathfrak{Z}^{+}(C_{n})\cong Q_{n}-V(K_{1,n})$. Since $Q_{n}-V(K_{1,n})$ has $12^{n-1}-$ 1 vertices of even cardinality and $2^{n-1}-n$ vertices of odd cardinality, $\mathfrak{Z}^{+}(C_{n})$ does not have a Hamilton path for $n\geq 3$.
Example 4.4.
Let $q\geq 1$. Since $K_{1,q}$ is a tree, $\operatorname{Z}_{+}(K_{1,q})=1=\overline{\operatorname{Z}}_{+}(K_{1,q})$ and $\mathfrak{Z}^{+}(K_{1,q})\cong Q_{q+1}-v$ for $v\in V(Q_{q+1})$. Thus $\mathfrak{Z}^{+}(K_{1,q})$ is not unique for $q\geq 3$ and $\mathfrak{Z}^{+}(K_{1,q})$ does have a Hamilton path.
For $K_{2,2}$, every set of two vertices is a PSD forcing set and no one vertex can force. For $q\geq 3$, the minimal PSD forcing sets of $K_{2,q}$ are $A$, $B$ and $\{a_{i},b_{j}\}$ where $a_{i}\in A$ and $b_{j}\in B$. For $p\geq 3$, a set $S\subseteq V(K_{a,b})$ is a PSD forcing set if and only if $A\subseteq S$, $B\subseteq S$, ($|S\cap A|=a-1$ and $|S\cap B|\geq 1$) or ($|S\cap B|=b-1$ and $|S\cap A|\geq 1$). Thus $\operatorname{Z}_{+}(K_{p,q})=p$ and $\overline{\operatorname{Z}}_{+}(K_{p,q})=q$ for $q\geq p\geq 2$.
The next result follows from Theorem 3.4 together with Remark 2.12, since $K_{2,q}$ has the same minimal sets for PSD forcing and domination (cf. Examples 4.4 and 3.2).
Corollary 4.5.
For $q\geq 3$, $\mathfrak{Z}^{+}(K_{2,q})$ has a Hamilton path if and only if $q$ is odd (and $\mathfrak{Z}^{+}(K_{2,q})$ does not have a Hamilton cycle).
Proposition 4.6.
For $q\geq 2$, $\mathfrak{Z}^{+}(K_{2,q})$ is unique.
Proof.
As usual, $B$ is the partite set with $q$ vertices and $A$ is the partite set with $p=2$ vertices. The cases $q=2,3$ are straightforward to verify computationally; see [11]. So assume $q\geq 4$. The minimal PSD forcing sets of $K_{2,q}$ are $A$, $B$ and $\{a_{i},b_{j}\}$ where $a_{i}\in A$ and $b_{j}\in B$. Suppose $G$ has no isolated vertices and $\mathfrak{Z}^{+}(G)\cong\mathfrak{Z}^{+}(K_{2,q})$. By Theorem 2.23, $\operatorname{Z}_{+}(G)=2$ and we can relabel the vertices of $G$ so that the PSD forcing sets are the same as for $K_{2,q}$. Recall that $\operatorname{Z}_{+}(G)=2$ implies $G$ cannot contain a $K_{4}$ minor [14]. We consider two cases, depending on whether or not the vertices in $A$ are adjacent.
First assume $G[A]=2K_{1}$. If $N_{G}(a_{1})\neq B$, then $N_{G}(a_{1})$ is a PSD forcing set, contradicting that $B$ is a minimal PSD forcing set. So $N_{G}(a_{1})=B=N_{G}(a_{2})$. Let $C$ be a connected component of $G[B]$. Since $a_{1}$ and $a_{2}$ are each adjacent to every vertex of $B$ and $A$ is a PSD forcing set, $|C|=1$. Thus $G\cong K_{2,q}$.
Now assume $G[A]=K_{2}$. We show first that for any vertex $b_{i}\in B$, $|N_{G}(b_{i})\cap A|=1\implies B\subseteq N_{G}[b_{i}]$. Suppose $|N_{G}(b_{i})\cap A|=1$. The one neighbor of $b_{i}$ in $A$, say $a_{1}$, can be forced blue by $b_{i}$. Then $a_{1}$ forces $a_{2}$ so $A$ is blue. Thus $N_{G}[b_{i}]\cap B$ is a PSD forcing set and the miminality of $B$ as a PSD forcing set implies $B\subseteq N_{G}[b_{i}]$. Since $B$ is a PSD forcing set, there exists a vertex $b_{k}$ such that $N_{G}(b_{k})\cap A=\{a_{1}\}$ (possibly after relabeling $a_{1}$ and $a_{2}$). Thus $N_{G}[b_{k}]=B\cup\{a_{1}\}$. Since a leaf and its one neighbor cannot both be in a minimal PSD forcing set, and for every pair of vertices there exists a minimal PSD forcing set that contains those vertices, $\delta(G)\geq 2$. Thus there exists a vertex $b_{\ell}\in B$ such that $b_{\ell}\in N_{G}(a_{2})$ (necessarily $\ell\neq k$ since $N_{G}(b_{k})\cap A=\{a_{1}\}$). Let $b_{i}\in B$ and $i\neq k,\ell$. If $A\subseteq N_{G}(b_{i})$, then we see that $G[\{a_{1},a_{2},b_{k},b_{\ell},b_{i}\}]$ would have a $K_{4}$ minor by contracting the edge $b_{k}b_{\ell}$. Since this is not allowed, $A\not\subseteq N_{G}(b_{i})$. If $A\cap N_{G}(b_{i})=\emptyset$, then $B\setminus\{b_{i}\}$ would be a PSD forcing set. So $A\cap N_{G}(b_{i})\neq\emptyset$. This implies that $|N_{G}(b_{i})\cap A|=1$, which implies $B\subset N_{G}[b_{i}]$. This is a contradiction because then contracting edge $a_{1}a_{2}$ gives a $K_{4}$ minor. Thus $G[A]\not\cong K_{2}$. ∎
Table 4.1 shows the number of (nonisomorphic) graphs without isolated vertices of order at most eight that have unique $\operatorname{Z}_{+}$-TAR graphs (this data was computed in [11]).
# vertices in $G$ 2 3 4 5 6 7 8
# graphs with unique $\mathfrak{Z}^{+}(G)$ 1 2 3 10 48 398 6798
# graphs with no isolated vertices 1 2 7 23 122 888 11302
ratio (# unique/# no isolated) 1 1 0.4286 0.4348 0.3934 0.4482 0.6015
Table 4.1: Number of graphs having a unique PSD TAR graph for graphs of small order
We do not have an example of a base graph $G$ such that $\mathfrak{Z}^{+}(G)$ has a Hamilton cycle.
4.2 Connectedness properties of the PSD TAR graph
The least $k_{0}$ such that $\mathfrak{Z}^{+}_{k}(G)$ is connected for all $k\geq k_{0}$ is denoted by $z^{+}_{0}(G)$ and the least $k$ such that $\mathfrak{Z}^{+}_{k}(G)$ is connected is denoted by $\underline{z^{+}_{0}}(G)$. As usual, there are many example of graphs having $\overline{\operatorname{Z}}_{+}(G)+1=z^{+}_{0}(G)$ and $\underline{z^{+}_{0}}(G)=z^{+}_{0}(G)$, such as complete graphs, trees, and cycles. Here we provide examples of graphs $G$ with $\overline{\operatorname{Z}}_{+}(G)+1<z^{+}_{0}(G)$ and $\underline{z^{+}_{0}}(G)<z^{+}_{0}(G)$
Proposition 4.7.
For $q\geq p\geq 2$, $\operatorname{Z}_{+}(K_{p,q})=p$ and $\overline{\operatorname{Z}}_{+}(K_{p,q})=q$. If $p\geq 3$ then $z^{+}_{0}(K_{p,q})=p+q-2$. Thus for $q=p\geq 4$, $z^{+}_{0}(K_{p,p})=2p-2=\overline{\operatorname{Z}}_{+}(K_{p,p})+p-2>\overline{\operatorname{Z}}_{+}(K_{p,p})+1$. For $q\geq 4$, $\underline{z^{+}_{0}}(K_{2,q})=3$ and $z^{+}_{0}(K_{2,q})=q+1$, so $z^{+}_{0}(K_{2,q})=q+1=\underline{z^{+}_{0}}(K_{2,q})+q-2>\underline{z^{+}_{0}}(K_{2,q})$.
Proof.
For $p\geq 3$, a set $S\subseteq V(K_{p,q})$ is a PSD forcing set if and only if $A\subseteq S$, $B\subseteq S$, ($|S\cap A|=p-1$ and $|S\cap B|\geq 1$) or ($|S\cap B|=q-1$ and $|S\cap A|\geq 1$). Thus $\operatorname{Z}_{+}(K_{p,p})=p=\overline{\operatorname{Z}}_{+}(K_{p,p})$ and $\mathfrak{Z}^{+}(K_{p,p})$ is disconnected until a set $S$ can contain at least $p-1$ vertices from each part, i.e. $|S|\geq 2p-2$. Furthermore, ${\mathfrak{Z}^{+}}_{k}(K_{p,p})$ is connected for $k\geq 2p-2$.
For $q\geq 4$, a set $S\subseteq V(K_{2,q})$ is a PSD forcing set if and only if $A\subseteq S$, $B\subseteq S$, or ($|S\cap A|=1$ and $|S\cap B|\geq 1$). Thus the minimal PSD forcing sets are $A,B$, and $S$ such that $|S\cap A|=1$ and $|S\cap B|=1$. Thus ${\mathfrak{Z}^{+}}_{2}(K_{2,q})$ is disconnected, ${\mathfrak{Z}^{+}}_{3}(K_{2,q})$ is connected, ${\mathfrak{Z}^{+}}_{q}(K_{2,q})$ is disconnected, and ${\mathfrak{Z}^{+}}_{k}(K_{2,q})$ is connected for $k\geq q+1$. ∎
5 TAR reconfiguration for skew forcing
In this section we define the skew forcing TAR graph, apply results of Section 2, and obtain additional results for skew TAR graphs, including results about uniqueness, existence or nonexistence of Hamilton paths and cycles, and connectedness. Denote the skew TAR graph of a base graph $G$ by $\mathfrak{Z}^{-}(G)$ and deonte the upper skew forcing number of $G$ by $\overline{\operatorname{Z}}_{-}(G)$.
Since skew forcing is a robust $X$-set parameter (but not an original $X$-set parameter), all the results that are true for robust $X$-set parameters and $\mathfrak{X}(G)$ apply to skew forcing and the skew TAR graph. Here we restate only the main result for isomorphic skew TAR graphs.
Theorem 5.1.
Suppose base graphs $G$ and $G^{\prime}$ have no isolated vertices and $\mathfrak{Z}^{-}(G)\cong\mathfrak{Z}^{-}(G^{\prime})$. Then $G$ and $G^{\prime}$ have the same order and there is a relabeling of the vertices of $G^{\prime}$ such that $G$ and $G^{\prime}$ have exactly the same skew forcing sets.
5.1 Examples of skew TAR graphs, uniqueness, and Hamiltonicity
In this section, we exhibit skew TAR graphs of some families of base graphs, including complete graphs, complete multipartite graphs, odd cycles, and every graph $G$ with $\operatorname{Z}_{-}(G)=0$ (which includes even paths). We show that the skew TAR graph of a complete multipartite graph is unique and examine Hamiltonicity.
Example 5.2.
It is well known that $\operatorname{Z}_{-}(K_{n})=n-2$ and any set of $n-2$ vertices is a skew forcing set. Thus $\mathfrak{X}(K_{n})$ is isomorphic to the top three ‘levels’ of $Q_{n}$, i.e., all the subsets of $n-2$, $n-1$, and $n$ vertices. It is shown in the next theorem that $\mathfrak{Z}^{-}(K_{n})$ is unique. For $n\geq 4$, $\mathfrak{Z}^{-}(K_{n})$ does not have a Hamilton path because there are $n$ sets of cardinality $n-1$ and $1+\frac{n(n-1)}{2}$ sets of cardinality $n$ or $n-2$ and $21+\frac{n(n-1)}{2}>n+$ 2.
For $t\geq 2$ and $n_{i}\geq 1,i=1,\dots,t$, a complete multiparite graph is a graph whose vertex set can be partitioned into sets of orders $n_{1},\dots,n_{t}$ such that the edge set is every edge between two distinct partite sets; this is denoted by $K_{n_{1},\dots,n_{t}}$.For a graph $G$ of order $n$, it is known that $\operatorname{Z}_{-}(G)=n-2$ if and only if $G$ is a complete multipartite graph [25, Theorem 9.76]. Note that the next results includes $K_{n}\cong K_{1,1,\dots,1}$ as well as complete bipartite graphs.
Theorem 5.3.
The complete multiparite graph $K_{n_{1},\dots,n_{t}}$ has as its minimal skew forcing sets the sets of $n-2$ vertices where the two vertices omitted are in different partite sets. These are the vertices of the skew TAR graph that have cardinality $n-2$ (and degree two). Every set of $n-1$ vertices of $K_{n_{1},\dots,n_{t}}$ is a vertex of the skew TAR graph and has degree $n-n_{i}+1$ where the vertex omitted is in the $i$th partite set. The skew TAR graph of a complete multiparite graph is unique.
Proof.
Suppose $\mathfrak{Z}^{-}(G)\cong\mathfrak{Z}^{-}(K_{n_{1},\dots,n_{t}})$ with $t\geq 2$ and $n_{i}\geq 1$, and let $n=n_{1}+\dots+n_{t}$. Then $|V(G)|=n$ and $\operatorname{Z}_{-}(G)=n-2$. Relabel the vertices of $G$ so that the skew forcing sets of $G$ are the same as the skew forcing sets of $K_{n_{1},\dots,n_{t}}$ (by Theorem 5.1); necessarily $V(G)=V(K_{n_{1},\dots,n_{t}})$. Denote the partite sets of $K_{n_{1},\dots,n_{t}}$ by $Y_{i}$ where $|Y_{i}|=n_{i}$. A set $S$ of $n-2$ vertices of $K_{n_{1},\dots,n_{t}}$ is a skew forcing set if and only if the two vertices not in $S$ are not in the same partite set. Consider a set $T=V(G)\setminus\{y_{i}\}$ of $n-1$ vertices where $y_{i}\in Y_{i}$. Removing a vertex $v$ of $T$ results in a skew forcing set $K_{n_{1},\dots,n_{t}}$ (and thus of $G$) if and only if $v\not\in Y_{i}$. Of course, $T\cup\{y_{i}\}$ is also a skew forcing set. Thus $\deg_{\mathfrak{Z}^{-}(G)}(T)=n-n_{i}+1$. We can recover $n_{1},\dots,n_{t}$ from the degrees of the vertices of $\mathfrak{Z}^{-}(G)$ that contain $n-1$ elements, because omitting a vertex from $Y_{i}$ contributes $n_{i}$ vertices of degree $n-n_{i}+1$. ∎
[…]
Observation 5.4.
The following are equivalent:
  1. $(1)$
    $\operatorname{Z}_{-}(G)=0$.
  2. $(2)$
    $\emptyset$ is a skew forcing set.
  3. $(3)$
    $\mathfrak{Z}^{-}(G)$ is a hypercube.
As noted in Remark 2.45, a hypercube has a Hamilton cycle.
The Leaf-Stripping Algorithm (reproduced below and implemented in [11]) can be used to determine whether $\operatorname{Z}_{-}(G)=0$, because $\operatorname{Z}_{-}(G)=0$ if and only if the algorithm returns the empty set [25, Algorithm 9.79, Theorem 9.80].
Algorithm 5.5 (Leaf-Stripping Algorithm).
[26, 31]
Input: Graph $G$.
Output: Graph $\hat{G}$ with $\delta(\hat{G})\neq 1$, or $\hat{G}=\emptyset$.
$\hat{G}:=G$.
While $\hat{G}$ has a leaf $u$ with neighbor $v$:
$\hat{G}:=\hat{G}[V(\hat{G})\setminus\{u,v\}]$.
Return $\hat{G}$.
Next we present examples of graphs in which any one vertex is a skew forcing set, and again nonisomorphic base graphs have isomorphic skew TAR graphs. It is well known that $\delta(G)-1\leq\operatorname{Z}_{-}(G)$ [25, Remark 9.63]. For $r\geq 2,s\geq 3$, the flower with $r$ petals of size $s$ (or the $(r,s)$-generalized friendship graph), denoted by $F(r,s)$, is a union of $r$ copies of $C_{s}$ that share one common vertex. Note that $F(r,s)$ has $(s-1)r+1$ vertices and $sr$ edges. The center vertex of a flower is the vertex that has degree at least four.
Proposition 5.6.
The following graphs $G$ have the property that any one vertex is a skew forcing set but the empty set is not a skew forcing set, so $\mathfrak{Z}^{-}(G)\cong Q_{n}-v$ for any $v\in V(Q_{n})$:
  1. $(1)$
    An odd cycle.
  2. $(2)$
    A flower $F(r,2k+1)$ for $r\geq 2,k\geq 1$.
As noted in Remark 2.45, a graph $Q_{n}-v$ for $v\in V(Q_{n})$ has a Hamilton path (but not a Hamilton cycle).
Proof.
In all cases, $\delta(G)=2$, so $\operatorname{Z}_{-}(G)\geq 1$.
(1): In an odd cycle, coloring any one vertex blue leaves an even path, which can then force itself (it is well known that $\operatorname{Z}_{-}(P_{2k})=0$).
(2): If the center vertex of is blue, the white vertices in each petal forms an even path, which can force itself. So suppose a degree-two vertex $u$ is blue. Since each petal has an odd number of vertices, there is a path with an even number of vertices from a white neighbor $w$ of $u$ to the center vertex $c$ (including $w$ and $c$). Using white vertex forcing along this path, $c$ can be forced. ∎
Returning to the question of uniqueness of skew TAR graphs, Table 5.1 shows the number of (nonismorphic) graphs without isolated vertices of order at most eight that have unique $\operatorname{Z}_{-}$-TAR graphs (this data was computed in [11]).
# vertices in $G$ 2 3 4 5 6 7 8
# graphs with unique $\mathfrak{Z}^{-}(G)$ 1 2 4 7 27 179 3026
# graphs with no isolated vertices 1 2 7 23 122 888 11302
ratio (# unique/# no isolated) 1 1 0.5714 0.3043 0.2213 0.2016 0.2677
Table 5.1: Number of graphs having a unique skew TAR graph for graphs of small order
For $n=2,3$ and $4$, the number of nonisomorphic graphs $G$ with unique $\mathfrak{Z}^{-}(G)$ is the same as the number of complete multipartite graphs of order $n$ (which is one less than the number of partitions of $n$). But beginning with $n=5$ there exists at least one additional base graph $G$ of each order that has a unique skew TAR graph. For $n=5$ there is only one, the Full House graph [11]; see Example 5.7.
222233330044441111
Figure 5.1: The Full House graph
Example 5.7.
The Full House graph, shown in Figure 5.1, has minimal skew forcing sets $\{3\},\{4\},\{0,1,2\}$.
5.2 Connectedness properties of the skew forcing TAR graph
In this section we discuss connectedness properties of the skew forcing TAR graph. The least $k_{0}$ such that $\mathfrak{Z}^{-}_{k}(G)$ is connected for all $k\geq k_{0}$ is denoted by $z^{-}_{0}(G)$ and the least $k$ such that $\mathfrak{Z}^{-}_{k}(G)$ is connected is denoted by $\underline{z^{-}_{0}}(G)$. We begin with a simple observation.
Observation 5.8.
Suppose the empty set is a skew forcing set of a graph $G$ of order $n$, which implies $\mathfrak{Z}^{-}(G)\cong Q_{n}$ and thus $\mathfrak{Z}^{-}_{k}(G)$ is connected for every $k\geq 0$. Therefore $z^{-}_{0}(G)=\underline{z^{-}_{0}}(G)=\operatorname{Z}_{-}(G)=0$ (cf. Proposition 2.34).
Many well known graphs have $z^{-}_{0}(G)=\overline{\operatorname{Z}}_{-}(G)+1$ and $\underline{z^{-}_{0}}(G)=z^{-}_{0}(G)$. Any graph for which $\operatorname{Z}_{-}(G)=0$ satisfies $\underline{z^{-}_{0}}(G)=z^{-}_{0}(G)$. It follows from results in Section 5.1 that the graphs $K_{n_{1},\dots,n_{t}},C_{n}$, and $F_{r}$ all have both these properties. We show that the family of graphs $H(r)$ presented in [7] to show that $z_{0}(G)$ can exceed the lower bound $\overline{\operatorname{Z}}(G)+1$ by an arbitrary amount also shows that $z^{-}_{0}(G)$ can exceed the lower bound $\overline{\operatorname{Z}}_{-}(G)+1$ by an arbitrary amount, although the values are different. We also apply the twin vertex method (Proposition 2.42) to construct a family of graphs that have $\underline{z_{0}}(G)$ strictly less than $z_{0}(G)$. Recall the graph $H(r)$ (see Section 3.3) is constructed from two copies of $K_{r+2}$ by adding a matching between $r$ pairs of vertices.
Proposition 5.9.
For $r\geq 2$, $\operatorname{Z}_{-}(H(r))=r$, $\overline{\operatorname{Z}}_{-}(H(r))=r+1$, and $z^{-}_{0}(H(r))=2r=\overline{\operatorname{Z}}_{-}(H(r))+r-1$.
Proof.
Observe that $\delta(H(r))=r+1$. Let the vertices of degree $r+1$ be denoted by $x,y,x^{\prime},y^{\prime}$ where $x$ and $y$ are adjacent, as are $x^{\prime}$ and $y^{\prime}$. Let $U=\{u_{1},\dots,u_{r}\}$ be the remaining neighbors of $x$, $U^{\prime}=\{u^{\prime}_{1},\dots,u^{\prime}_{r}\}$ be the remaining neighbors of $x^{\prime}$, $V=U\cup\{x,y\}$, and $V^{\prime}=U^{\prime}\cup\{x^{\prime},y^{\prime}\}$. Then $U$ is a skew forcing set, so $r=\delta(H(r))-1\leq\operatorname{Z}_{-}(H(r))\leq r.$ In fact, any set of $r$ vertices in $V$ that contains at most one of $x$ or $y$ is a skew forcing set (and similarly for $V^{\prime}$). A set $S$ of $r$ vertices in $V$ that contains both $x$ and $y$ is not a skew forcing set, but adding a neighbor of $v\in(V\setminus S)$ produces a minimal skew forcing set of $r+1$ vertices. No set of $r+2$ or more vertices of $H(r)$ is minimal, so $\overline{\operatorname{Z}}_{-}(H(r))=r+1$. We partition the minimal skew forcing sets into those that have at least $r$ vertices in $V$ and those that have at least $r$ vertices in $V^{\prime}$. Then by Proposition 2.38, $z^{-}_{0}(H(r))\geq(r+1)+(r+1)-2=2r$. Since $|S\cup T|\leq 2r$ for any minimal skew forcing sets $S$ and $T$, $z^{-}_{0}(H(r))=2r$. ∎
A family of graphs $H_{r}$ having $\underline{z_{0}}(H_{r})<z_{0}(H_{r})$ was constructed in [7] by using twins. This family also has the property that $\underline{z^{-}_{0}}(H_{r})<z^{-}_{0}(H_{r})$, but for skew forcing there are graphs with this property of smaller order, including the family obtained by creating twins of the degree two vertex in the Full House graph.
The next result is the skew analog of [7, Proposition 3.6], but follows from known results: Let $G$ be a graph with a set of $r$ independent twins. It is immediate that any skew forcing set of $G$ must contain at least $r-1$ of the vertices of $T$. The remaining two conditions for the twins property are established in the proof of Proposition 9.87 in [25].
Proposition 5.10.
Skew forcing has the twins property.
Now we construct the family $FH(r)$: Define $FH(1)=FH$ and construct $FH(r+1)$ from $FH(r)$ by adding another independent twin of a degree-two vertex; $FH(2)$ is shown in Figure 5.2
Figure 5.2: A graph $FH(2)$ satisfying $z^{-}_{0}(FH(2))>\underline{z^{-}_{0}}(FH(2))$
Proposition 5.11.
For $r\geq 1$, $\operatorname{Z}_{-}(FH(r))=r$, $\overline{\operatorname{Z}}_{-}(FH(r))=r+2$, $\underline{z^{-}_{0}}(FH(r))=r+1$, and $z^{-}_{0}(FH(r))=r+3$.
Proof.
The results for $r=1$ are in Example 5.7 and are computed in [11] for $r=2$. The first three equalities then follow from Proposition 2.42. Since $\overline{\operatorname{Z}}_{-}(FH(r))+1\leq z^{-}_{0}(FH(r))$, ${\mathfrak{Z}^{-}}_{r+3}(FH(r))$ is connected, and $|V(FH(r))|=r+4$, we have $z^{-}_{0}(FH(r))=r+3$. ∎
5.3 Skew irrelevant vertices
Recall that a vertex $v$ of a graph $G$ is skew irrelevant if it is not in any minimal skew forcing set of $G$. Irrelevant vertices play a key role in the automorphism group of an $X$-TAR graph. In this section we examine skew irrelvant vertices.
Observation 5.12.
If a vertex $v$ of $G$ can be skew forced by the empty set, then $v$ is skew-irrelevant.
We see in the Example 5.14 that the converse of Observation 5.12 is not true. However, it is true for trees.
Proposition 5.13.
Let $T$ be a forest. A vertex $v$ of $T$ is skew irrelevant if and only if it can be skew forced by the empty set.
Proof.
Observe that applying the Leaf-Stripping Algorithm 5.5 to a forest always returns a set (possibly empty) of isolated vertices. During any application of the algorithm, $\hat{T}$ is a forest after each iteration. Let $v\in V(T)$. Since a forest that has an edge must have at least two leaves, we can always choose to not have $v$ perform a force even if $v$ is a leaf of $\hat{T}$. So at the end of the algorithm either $v$ was forced by the empty set or $v\in V(\hat{T})$ and $V(\hat{T})$ is a minimum skew forcing set. ∎
Example 5.14.
Let $F(r)$ denote the graph constructed from $r$ copies of $K_{3}$ and one $K_{1,r}$, where each leaf of the $K_{i,r}$ is a vertex of one $K_{3}$. The graph $F(3)$ is shown in Figure 5.3. Then every minimal skew forcing set of $F(r)$ has $r-1$ vertices taken from distinct copies of $K_{3}$. Notice that the center vertex of the $K_{1,r}$ is irrelevant but no vertex can be forced by the empty set. The irrelevant vertex means that the skew TAR graph is a Cartesian product: Let $W$ be the set of skew forcing sets of $F(r)$ that do not contain the irrelevant vertex. Then $\mathfrak{Z}^{-}(F(r))\cong\left(\mathfrak{Z}^{-}(F(r))[W]\right)\Box P_{2}$.
Figure 5.3: The graph $F(3)$
6 TAR reconfiguration for vertex covers
In this section we define the vertex cover TAR graph, apply results of Section 2, and obtain additional results for such TAR graphs, including uniqueness, connectedness, and irrelevant vertex results. Recall that the vertex cover number of $G$ is denoted by $\tau(G)$, so the upper vertex cover number of $G$ is denoted by $\overline{\tau}(G)$. Denote vertex cover TAR graph or VC TAR graph of a base graph $G$ by $\mathfrak{C}(G)$. As with the other parameters, our focus here is on isomorphisms of TAR graphs, connectedness, existence of Hamilton cycles or paths, etc. There is a close relationship between vertex cover reconfiguration and independent set reconfiguration (see Section 8.3), and there is is extensive prior work on reconfiguration of both independent sets and vertex covers, including TAR reconfiguration; often the work on independent sets was done first. Most of this work is focused on complexity, approximation, and algorithms (see, for example, [24, 28, 29, 30]). However, we point out that some of these papers provide useful tools for the type of results we emphasize. For instance, results from [29] and [30] are used to establish equality in connectedness bounds for certain families of graphs in Propositions 6.8 and 8.22.
Since the vertex cover number is a robust $X$-set parameter (but not an original $X$-set parameter because $\tau(K_{1})=0$), all the results that are true for robust $X$-set parameters and $\mathfrak{X}(G)$ apply to vertex covers and the VC TAR graph. Here we restate only the main result for isomorphic VC TAR graphs.
Theorem 6.1.
Let $G$ and $G^{\prime}$ be graphs. If $\mathfrak{C}(G)\cong\mathfrak{C}(G^{\prime})$, then $G$ and $G^{\prime}$ have the same order and there is a relabeling of the vertices of $G^{\prime}$ such that $G$ and $G^{\prime}$ have exactly the same vertex covers.
The robust $X$-set parameter vertex cover number is the base graph complement parameter to the robust $Y$-set parameter independence number (in the sense that $S$ is independent if and only if $V(G)\setminus S$ is a vertex cover); more information about the base graph complement parameter relationship can be found in Section 8.2 and more information about independence number can be found in Section 8.3.
6.1 Examples of VC TAR graphs, uniqueness, and Hamiltonicity
In this section, we exhibit vertex cover TAR graphs of some families of base graphs, including complete graphs, complete bipartite graphs, and empty graphs, including examples with no Hamilton path, Hamilton path but no Hamilton cycle, and Hamilton cycle. We also show every vertex cover TAR graph is unique.
Example 6.2.
Every vertex cover for $K_{n}$ needs $n-1$ vertices. Thus $\tau(K_{n})=n-1=\overline{\tau}(K_{n})$, and $\mathfrak{C}(K_{n})\cong K_{1,n}$. Furthermore, $K_{n}$ does not have a Hamilton path for $n\geq 3$.
Example 6.3.
The two partite sets $A$ and $B$ of $K_{p,q}$ are minimal vertex covers and these are the only minimal vertex covers. Thus $\mathfrak{C}(K_{p,q})$ is the vertex sum of hypercubes $Q_{p}$ and $Q_{q}$, $V(K_{p,q})=A\,\sqcup\,B$ is a cut-vertex of $\mathfrak{C}(K_{p,q})$, $\tau(K_{p,q})=p$, $\overline{\tau}(K_{p,q})=q$, and $\tau_{0}(K_{p,q})=p+q$. Note that $\mathfrak{C}_{q}(K_{p,q})$ is disconnected, but if $p<q$, then $\underline{\tau_{0}}(K_{p,q})=p$. Furthermore, $\mathfrak{C}(K_{p,q})$ has a Hamilton path but not a Hamilton cycle.
Example 6.4.
A path on $n$ vertices has $\tau(P_{n})=\left\lfloor\frac{n}{2}\right\rfloor$. Label the vertices of $P_{n}$ in path order. Using the hypercube representation, a subset $S$ of $V(P_{n})$ is represented by the sequence $(s_{1},\dots,s_{n})$ where $s_{i}=1$ if $i\in S$ and $s_{i}=0$ if $i\not\in S$. A set $S$ is a vertex cover for $P_{n}$ if and only if the sequence for $S$ has no consecutive zeros. As is standard in the hypercube representation, two vertices of $\mathfrak{C}(P_{n})$ are adjacent if and only if they differ in exactly one digit. In particular, $P_{4}$ has vertex cover sets $1111$, $1110$, $1101$, $1011$, $0111$, $1010$, $0101$ and $0110$. A Hamilton cycle for $\mathfrak{C}(P_{4})$ is described by the sequence
$(1111,1011,1010,1110,0110,0111,0101,1101,1111)$
Example 6.5.
Since the vertex cover number is a robust $X$-set parameter and $\tau(K_{1})=0$, $\mathfrak{C}(\overline{K_{n}})\cong K_{2}\Box\dots\Box K_{2}\cong Q_{n}$ for $n\geq 1$ by Proposition 2.8. Thus $\mathfrak{C}(\overline{K_{n}})$ has a Hamilton cycle.
Proposition 6.6.
Every vertex cover TAR graph is unique.
Proof.
Let $G$ be a graph. For $u,w\in V(G)$, the set $V(G)\setminus\{u,w\}$ is a vertex cover if and only if $uw\not\in E(G)$. Thus the vertex cover TAR graph $\mathfrak{C}(G)$ determines the base graph. Suppose $\mathfrak{C}(G)\cong\mathfrak{C}(G^{\prime})$. By Theorem 6.1, we may relabel the vertices of $G^{\prime}$ to obtain $G^{\prime\prime}$ such that $\mathfrak{C}(G)$ and $\mathfrak{C}(G^{\prime\prime})$ have the same vertex covers. Thus $G$ and $G^{\prime\prime}$ are the same graph, and $G=G^{\prime\prime}\cong G^{\prime}$. So the VC TAR graph is unique. ∎
6.2 Connectedness properties of the VC TAR graph
The least $k_{0}$ such that $\mathfrak{C}_{k}(G)$ is connected for all $k\geq k_{0}$ is denoted by $\tau_{0}(G)$ and the least $k$ such that $\mathfrak{C}_{k}(G)$ is connected is denoted by $\underline{\tau_{0}}(G)$. Although complexity and algorithms are main interests of Ito, Nooka, and Zhou in [29], they also establish equality in a lower bound for $\tau_{0}(G)$ for certain families of graphs $G$.
Remark 6.7.
Let $G$ be a graph and let $C_{0}$ and $C_{t}$ be two vertex covers of $G$. As defined [29], the minmax vertex cover reconfigutation problem is to determine the least index $k$ such that there is a path between $C_{0}$ and $C_{t}$ in $\mathfrak{C}_{k}(G)$. The minmax vertex cover reconfigutation problem is one of the main problems studied in [29]. When maximized over all pairs $C_{0},C_{t}$, this is equivalent to determining $\tau_{0}(G)$.
There are many examples of graphs that having $\tau_{0}(G)>\overline{\tau}(G)+1$, including complete bipartite graphs; see, e.g., Example 6.3 with $p\geq 2$ or $p=1,q\geq 3$. As shown in that example, $\tau_{0}(K_{p,q})=\overline{\tau}(K_{p,q})+\tau(K_{p,q})$, so $\underline{\tau_{0}}(K_{p,q})=p<p+q=\tau_{0}(K_{p,q})$. The next result provides examples where $\tau_{0}(G)=\overline{\tau}(G)+1$. An even–hole–free graph is a graph which contains no induced even cycles. The class of even–hole–free graphs includes trees, chordal graphs and interval graphs (see [13, Section 5.5] for definitions of chordal graphs and interval graphs).
Proposition 6.8.
Let $G$ be an even–hole–free graph. Then $\tau_{0}(G)=\overline{\tau}+1$.
Proof.
Lemma 2 in [29] uses an analogous result for independent sets in [30] (see Theorem 8.21) to show that if $S_{1}$ and $S_{2}$ are vertex covers of a graph $G$, then there exists a path in $\mathfrak{C}_{k}(G)$ for $k=\max(|S_{1}|,|S_{2}|)+1$. Let $\ell=\overline{\tau}(G)$ and let $S_{1},S_{2}\in V(\mathfrak{C}_{\ell+1}(G))$. Then there exists minimal vertex covers $M_{1}$ and $M_{2}$ such that $M_{i}\subseteq S_{i},i=1,2$. Since $|M_{i}|\leq\ell,i=1,2$, there exists a path in $\mathfrak{C}_{\ell+1}(G)$ from $M_{1}$ to $M_{2}$. Then starting at $S_{1}$ and removing the vertices in $S_{1}\setminus M_{1}$ one at a time, traversing the path from $M_{1}$ to $M_{2}$, and then adding the vertices in $S_{2}\setminus M_{2}$ one at a time gives a path from $S_{1}$ to $S_{2}$. Since $\mathfrak{C}_{\ell+1}(G)$ is connected, $\tau_{0}(G)=\overline{\tau}(G)+1$ by Propositions 2.36(2) and 2.34(3). ∎
6.3 VC-irrelevant vertices
A vertex $v$ is VC-irrelevant if it is not in any minimal vertex cover.
Proposition 6.9.
Let $G$ be a graph and $v\in V(G)$. Then $V$ is $VC$-irrelevant if and only if $v$ is an isolated vertex.
Proof.
An isolated vertex is not in any minimal vertex cover. Now suppose $u$ is incident with an edge. A minimal vertex cover containing $u$ can be constructed by starting with $S=\{u\}$ and repeatedly adding one vertex at a time, choosing a vertex $w$ that is not an endpoint of any edge that has already been covered, until all edges are covered. ∎
7 Connected domination
Connected domination is a parameter that has been studied in the literature [9, 35] and is an example of a super $X$-set parameter that is not robust, and in fact, not a connected $X$-set parameter, because it does not satisfy the $(n-1)$-set axiom. Furthermore, graphs of different orders can have the same connected domination TAR graphs. For a connected graph $G$, a connected dominating set is a (standard) dominating set $S$ of a graph $G$ such that $G[S]$ is connected, and the connected domination number, $\gamma_{c}(G)$, is the minimum cardinality of a connected dominating set. As noted in [9], every superset of a connected dominating set is a connected dominating set, because if $S$ is a connected dominating set then every vertex not in $S$ is a neighbor of a vertex in $S$. Thus connected domination is a super $X$-set parameter, and the connected domination TAR reconfiguration graph of a base graph $G$ is denoted by $\mathfrak{D}^{c}(G)$. However, as seen in the next example, connected domination does not satisfy the $(n-1)$-set axiom, so it is not a connected $X$-set parameter (and is not robust). The next example shows it is possible to have $\mathfrak{D}^{c}(G)\cong Q_{r}$ for some $r<|V(G)|$. Example 7.2 shows that base graphs of different orders can have the same connected domination TAR graph.
Example 7.1.
Consider the star $K_{1,n-1}$. For $n\geq 3$, every connected dominating set contains the center vertex. Thus the set of $n-1$ leaf vertices is not a connected dominating set. Furthermore, $\mathfrak{D}^{c}(K_{1,n-1})\cong Q_{n-1}$.
For $r,s,t\geq 2$, a double-broom $DB_{r}(s,t)$ is a tree obtained by attaching $s$ leaves to one end vertex of a path $P_{r}$ and $t$ leaves to the other end vertex of the path.
Example 7.2.
Observe that every connected dominating set of $DB(r,s,t)$ must contain the $r$ vertices of the path. Furthermore, $\mathfrak{D}^{c}(DB_{r}(s,t))\cong Q_{s+t}$. Thus base graphs of different orders can have isomorphic connected domination TAR graphs.
8 $Y$-set parameters (subsets and maximal sets)
The original $X$-set parameters were strongly motivated by parameters related to standard zero forcing and domination. As such, the abstraction of these parameters naturally led to the Superset axiom. Another natural family of cohesive parameters is obtained by replacing the Superset axiom with a “Subset axiom.” Many of the results for super $X$-set parameters, including connectedness results, can be naturally adapted to sub $Y$-set parameters. The main isomorphism theorem, Theorem 2.23, is extended to certain sub $Y$-set parameters through two complementation techniques, described in Sections 8.2 and 8.5. This allows us to apply results to additional parameters such as independence number, irredundance, and failed zero forcing.
Definition 8.1.
A sub $Y$-set parameter is a cohesive parameter $Y$ such that $Y(G)$ is defined to be the maximum cardinality of a $Y$-set of $G$ where the $Y$-sets of $G$ satisfy the following condition:
  1. (I)
    (Subset) If $T$ is a $Y$-set of $G$ and $T^{\prime}\subset T$, then $T^{\prime}$ is a $Y$-set of $G$.
Remark 8.2.
Let $Y$ be a sub $Y$-set parameter and let $G$ be a graph. Recall that an abstract simplicial complex $\mathcal{C}$ is a collection of finite sets such that if $A\in\mathcal{C}$ and $B\subseteq A$, then $B\in\mathcal{C}$ (see, for example, [36, Definition 4.1.2]). Furthermore, the set of vertices of $\mathcal{C}$ is $V(\mathcal{C})=\cup_{A\in\mathcal{C}}A$. In the setting of matroids an abstract simplicial complex is referred to as an independence system. The $Y$-sets of $G$ form an abstract simplicial complex. Note that $V(\mathcal{C})\subseteq V(G)$ but these sets need not be equal (See Example 8.34).
When $Y$ is a sub $Y$-set parameter, the $Y$-TAR graph of a base graph $G$ is denoted by $\mathfrak{Y}(G)$ and the lower $Y$ number, denoted by $\underline{Y}(G)$, is the minimum cardinality of a maximal $Y$-set.
8.1 Connectedness
The study of $k$-TAR reconfiguration graphs in the setting of sub $Y$-set parameters requires minor modifications to Definition 2.33.
Definition 8.3.
Suppose $Y$ is a sub $Y$-set parameter. The $k$-token addition and removal (TAR) reconfiguration graph for $Y$, denoted by $\mathfrak{Y}_{k}(G)$, is the subgraph of $\mathfrak{Y}(G)$ induced by the set of all $Y$-sets of cardinality at least $k$.
The greatest $k_{0}$ such that $\mathfrak{Y}_{k}(G)$ is connected for all $k\leq k_{0}$ is denoted by $y_{0}(G)$, and the greatest $k$ such that $\mathfrak{Y}_{k}(G)$ is connected is denoted by $\overline{y_{0}}(G)$.
The next result is the $Y$-set parameter version of Proposition 2.34. The proofs of (1), (2), and (4) are analogous; we prove (3) here since it needs more significant adjustment.
Proposition 8.4.
Let $Y$ be a sub $Y$-set parameter and let $G$ be a graph of order $n$.
  1. $(1)$
    Then $Y(G)\geq\overline{y_{0}}(G)\geq y_{0}(G)$.
  2. $(2)$
    If $G$ has only one maximal $Y$-set, then $Y(G)=\overline{y_{0}}(G)=y_{0}(G)$.
  3. $(3)$
    If $G$ has more than one maximal $Y$-set, then $\underline{Y}(G)-1\geq y_{0}(G)\geq\max\{\underline{Y}(G)+Y(G)-n,0\}$.
  4. $(4)$
    If $G$ has more than one maximum $Y$-set, then $Y(G)-1\geq\overline{y_{0}}(G)$.
Proof.
(1): This relationship follows immediately from the definitions.
(2): Taking away one vertex at a time from one maximal $Y$-set does not disconnect the graph, therefore $\mathfrak{Y}_{k}(G)$ is always connected for $k\leq Y(G)$ when there is only one maximal $Y$-set.
(3): Let $\hat{T}\subseteq V(G)$ be a maximal $Y$-set with $|\hat{T}|=\underline{Y}(G)$. Then $\hat{T}$ is an isolated vertex of $\mathfrak{Y}_{\underline{Y}(G)}(G)$ (because we can’t remove a vertex, and adding a vertex results in a set that is not a $Y$-set). Thus $\underline{Y}(G)-1\leq y_{0}(G)$. It is immediate from the definition of $y_{0}(G)$ that $y_{0}(G)\geq 0$. Suppose that $\underline{Y}(G)+Y(G)-n\geq 0$ and let $k_{0}=\underline{Y}(G)+Y(G)-n$. Let $T\subseteq V(G)$ be a maximal $Y$-set of $G$ and let $T^{\prime}\subseteq V(G)$ be a maximum $Y$-set of $G$. To ensure $\mathfrak{Y}_{k}(G)$ is connected for all $k\leq k_{0}$, it is sufficient to show that every such pair of vertices $T$ and $T^{\prime}$ is connected in $\mathfrak{Y}_{k_{0}}(G)$. Define $T^{\prime\prime}=T\cap T^{\prime}$ and observe that $|T^{\prime\prime}|\geq k_{o}$. Then each of $T$ and $T^{\prime}$ is connected by a path to $T^{\prime\prime}$ by removing one vertex at a time. Thus $y_{0}(G)\geq\underline{Y}(G)+Y(G)-n$.
(4): Each minimum $Y$-set is an isolated vertex in $\mathfrak{Y}_{Y(G)}(G)$.
It is often easy to find examples of graphs for which $\underline{Y}(G)-1=y_{0}(G)$ and $\overline{y_{0}}(G)=y_{0}(G)$. The next result, which is the $Y$-set parameter version of Corollary 2.35, provides some such examples.
Corollary 8.5.
Let $Y$ be a sub $Y$-set parameter. If $G$ has more than one maximal $Y$-set and $Y(G)=n-1$, then $y_{0}(G)=\underline{Y}(G)-1$.
The next result is the $Y$-set parameter version of Proposition 2.36; the proof is analogous.
Proposition 8.6.
Let $Y$ be a sub $X$-set parameter and let $G$ be a graph of order $n$.
  1. $(1)$
    If for every pair of maximal $X$-sets $M_{1}$ and $M_{2}$, there is a path between $M_{1}$ and $M_{2}$ in $\mathfrak{Y}_{k}(G)$, then $\mathfrak{Y}_{k}(G)$ is connected.
  2. $(2)$
    If $k\leq\underline{Y}(G)$ and $\mathfrak{Y}_{k}(G)$ is connected, then $y_{0}(G)\geq k$.
  3. $(3)$
    If $|M_{1}\cap M_{2}|\geq k$ for every pair of maximal $Y$-sets $M_{1}$ and $M_{2}$, then $y_{0}(G)\geq k$.
Proof.
Suppose first that for every pair of maximal $X$-sets $M_{1}$ and $M_{2}$, there is a path between $M_{1}$ and $M_{2}$ in $\mathfrak{Y}_{k}(G)$. Given two $Y$-sets $T_{1},T_{2}\in V(\mathfrak{Y}_{k}(G))$, each $T_{i}$ is contained in a maximal $Y$-set $M_{i}$. There are paths in $\mathfrak{Y}_{k}(G)$ from $T_{1}$ to $M_{1}$, $M_{1}$ to $M_{2}$, and $M_{2}$ to $T_{2}$, so $\mathfrak{Y}_{k}(G)$ is connected.
Now assume $k\leq\underline{Y}(G)$, $\mathfrak{Y}_{k}(G)$ is connected, and $\ell<k$. Since $k\geq\underline{Y}(G)$, $\mathfrak{Y}_{k}(G)$ contains every maximal $Y$-set and there is a path between every pair of maximal $Y$-sets in $\mathfrak{Y}_{k}(G)$, which is a subgraph of $\mathfrak{Y}_{\ell}(G)$. Thus $\mathfrak{Y}_{\ell}(G)$ is connected by (1).
If $|M_{1}\cap M_{2}|\geq k$ for every pair of minimal $X$-sets $M_{1}$ and $M_{2}$, then there is a path through $M_{1}\cap M_{2}$ in $\mathfrak{Y}_{k}(G)$ for every pair of maximal $Y$-sets $M_{1}$ and $M_{2}$ and $y_{0}(G)\geq k$. ∎
The next result extends Theorem 7 in [30] from independent sets to $Y$-sets, and the same proof remains valid (the proof of Proposition 2.44 is essentially a complemententary version of that proof in [30]).
Proposition 8.7.
Let $Y$ be a sub $Y$-set parameter, let $G$ be a graph, and let $T_{1}$ and $T_{2}$ be $Y$-sets of $G$ with $|T_{i}|=k,i=1,2$. Then there is a path between $T_{1}$ and $T_{2}$ in the $k$-TJ reconfiguration graph of $G$ if and only if there is a path between $T_{1}$ and $T_{2}$ in $\mathfrak{Y}_{k-1}(G)$.
8.2 Base graph complement parameters
A key to studying sub $Y$-set parameters is observing the following correspondence between sub $Y$-sets and super $X$-sets obtained by taking set complements in the base graph.
Definition 8.8.
(Base graph complement parameters) Let $Y$ be a sub $Y$-set parameter and let $G$ be a graph. We say that a subset $S\subseteq V(G)$ is an $X_{Y}$-set if and only if $V(G)\setminus S$ is a $Y$-set of $G$, and we define the parameter $X_{Y}(G)$ to be the minimum cardinality of an $X_{Y}$-set of $G$.
Let $X$ be a super $X$-set parameter. We say that a subset $T\subseteq V(G)$ is a $Y_{X}$-set if and only if $V(G)\setminus T$ is an $X$-set of $G$, and we define the parameter $Y_{X}(G)$ to be the maximum cardinality of a $Y_{X}$-set of $G$.
The parameters $Y$ and $X_{Y}$ (or $X$ and $Y_{X}$) are called base graph complement parameters. The $X_{Y}$-TAR and $Y_{X}$-TAR graphs of a base graph $G$ are denoted by ${\mathfrak{X}}_{Y}(G)$ and ${\mathfrak{Y}}_{X}(G)$
The following definition introduces a notion of robustness for sub $Y$-set parameters that is analogous to super $X$-set parameters via base graph complement parameters. A set of one vertex is called a singleton.
Definition 8.9.
A robust $Y$-set parameter is a sub $Y$-set parameter $Y$ such that $Y(G)$ and the $Y$-sets of $G$ satisfy the following conditions:
  1. (I)
    (Subset) If $T$ is a $Y$-set of $G$ and $T^{\prime}\subset T$, then $T^{\prime}$ is a $Y$-set of $G$.
  2. (II)
    (Singletons) If $G$ is a connected graph of order at least two, then every singleton is a $Y$-set.
  3. (III)
    (Component consistency) Let $G$ be a graph with connected components $G_{1},\ldots,G_{k}$. Then $T$ is a $Y$-set of $G$ if and only if $T\cap V(G_{i})$ is a $Y$-set of $G_{i}$ for $i=1,\dots,k$.
Remark 8.10.
The following statements are immediate from the definitions: If $Y$ is a sub (respectively, robust) $Y$-set parameter, then $X_{Y}$ is a super (respectively, robust) $X$-set parameter. If $X$ is a super (respectively, robust) $X$-set parameter, then $Y_{X}$ is a sub (respectively, robust) $Y$-set parameter. The $Y_{X_{Y}}$-sets of $G$ are the $Y$-sets of $G$ and $Y_{X_{Y}}(G)=Y(G)$ and analogous statements hold when $Y$ and $X$ are interchanged.
The definition of base graph complement parameters is illustrated in the next example.
Example 8.11.
Label the vertices of the star $K_{1,3}$ with $\{1,2,3,4\}$ where $1$ is the vertex of degree three. The minimal standard zero forcing sets of $K_{1,3}$ are $\{2,3\},\{2,4\},$ and $\{3,4\}$ and $\mathfrak{Z}(K_{1,3})$ is shown in Figure 8.1(a) with the part of the hypercube of all subsets that are not $\operatorname{Z}$-sets shown in gray. The maximal $Y_{\operatorname{Z}}$-sets of $K_{1,3}$ are $\{1,4\},\{1,3\},$ and $\{1,2\}$ and $\mathfrak{Y}_{\operatorname{Z}}(K_{1,3})$ is shown in Figure 8.1(b) with the part of the hypercube of all subsets that are not $Y_{\operatorname{Z}}$-sets shown in gray.
$\{1,2,3,4\}$$\{1,2,3\}$$\{1,2,4\}$$\{1,3,4\}$$\{2,3,4\}$$\{1,2\}$$\{1,3\}$$\{2,3\}$$\{1,4\}$$\{2,4\}$$\{3,4\}$$\{1\}$$\{2\}$$\{3\}$$\{4\}$$\emptyset$(a)$\emptyset$$\{4\}$$\{3\}$$\{2\}$$\{1\}$$\{3,4\}$$\{2,4\}$$\{2,3\}$$\{1,4\}$$\{1,3\}$$\{1,2\}$$\{2,3,4\}$$\{1,3,4\}$$\{1,2,4\}$$\{1,2,3\}$$\{1,2,3,4\}$(b)
Figure 8.1: The graphs $\mathfrak{Z}(K_{1,3})$ and $\mathfrak{Y}_{\operatorname{Z}}(K_{1,3})$, illustrating the isomorphism
Remark 8.12.
Let $Y$ be a sub $Y$-set parameter. For every graph $G$, $\mathfrak{Y}(G)\cong{\mathfrak{X}}_{Y}(G)$ by the isomorphism $T\to V(G)\setminus T$. This isomorphism is illustrated in Figure 8.1 by inverting the hypercube of all subsets of the base graph.
Next we translate the main isomorphism result to robust $Y$-set parameters.
Theorem 8.13.
Let $Y$ be a robust $Y$-set parameter and let $G$ and $G^{\prime}$ be base graphs such $\mathfrak{Y}(G)\cong\mathfrak{Y}(G^{\prime})$. If $Y(K_{1})=1$ or $G$ and $G^{\prime}$ have no isolated vertices, then $G$ and $G^{\prime}$ have the same order and there is a relabeling of the vertices of $G^{\prime}$ such that $G$ and $G^{\prime}$ have exactly the same $Y$-sets.
Proof.
Note that $\mathfrak{Y}(G)\cong{\mathfrak{X}}_{Y}(G))$ and $\mathfrak{Y}(G^{\prime})\cong{\mathfrak{X}}_{Y}(G^{\prime}))$ by the isomorphism $T\to V(G)\setminus T$, which implies ${\mathfrak{X}}_{Y}(G)\cong{\mathfrak{X}}_{Y}(G^{\prime})$. Since $Y(K_{1})=1$ or $G,G^{\prime}$ have no isolated vertices, we see that $X_{Y}(K_{1})=0$ or $G,G^{\prime}$ have no isolated vertices. Since $Y$ is a robust $Y$-set parameter, $X_{Y}$ is a robust $X$-set parameter. Thus $G$ and $G^{\prime}$ have the same order and there is a relabeling of the vertices of $G^{\prime}$ such that $G$ and $G^{\prime}$ have exactly the same $X_{Y}$-sets by Theorem 2.23. Therefore, $G$ and $G^{\prime}$ have exactly the same $Y$-sets. ∎
Recall that for a super $X$-set parameter, a vertex $v$ of $G$ is $X$-irrelevant if $v\not\in S$ for every minimal $X$-set $S$ of $G$. The next definition translates irrelevance from super $X$-set parameters to sub $Y$-set parameters using complementary parameters.
Definition 8.14.
Let $G$ be a graph and let $Y$ be a sub $Y$-set parameter. A vertex $v\in V(G)$ is $Y$-irrelevant if $v\in T$ for every maximal $Y$-set $T$ of $G$.
Observation 8.15.
Let $G$ be a graph and let $Y$ be a sub $Y$-set paramter. Then $v\in V(G)$ is $Y$-irrelevant if and only if $v$ is $X_{Y}$-irrelevant. Similarly for a super $X$-set parameter $X$, $v\in V(G)$ is $X$-irrelevant if and only if $v$ is $Y_{X}$-irrelevant.
Remark 8.16.
A connected $Y$-set parameter definition analogous to Definition 2.10 could also be given, allowing the application of these results to connected graphs when the $Y$-set parameter lacks component consistency.
8.3 Independent Sets
In this section we define the independence TAR graph, list some known results for such graphs, and apply results of about robust $Y$-set parameters. As with the other parameters, our focus here is on isomorphisms of TAR graphs, connectedness, existence of Hamilton cycles or paths, etc. There is extensive prior work on reconfiguration of independent sets, including TAR reconfiguration (see, for example [28, 30]); most of this work is focused on complexity, approximation, and algorithms. However, we point out that some of these papers provide useful tools for the type of results we emphasize. In particular, [30] contains structural results, including one that implies equality in a connectedness bound for certain families of graphs (see Proposition 8.22).
Observe that the independence number $\alpha$ is a robust $Y$-set parameter. Furthermore, the independence number is the base graph complement parameter of the vertex cover number. Denote the independence TAR graph of a base graph $G$ by $\mathfrak{I}(G)$. All the properties of a robust $Y$-set parameter apply to independence number. Here we state only the main theorem (noting that $\alpha(K_{1})=1$).
Theorem 8.17.
Let $G$ and $G^{\prime}$ be graphs. If $\mathfrak{I}(G)\cong\mathfrak{I}(G^{\prime})$, then $G$ and $G^{\prime}$ have the same order and there is a relabeling of the vertices of $G^{\prime}$ such that $G$ and $G^{\prime}$ have exactly the same independent sets.
Note that a vertex is $\alpha$-irrelevant if and only if it is an isolated vertex. Since independence number is complementary to vertex cover number, we have the next result.
Proposition 8.18.
Every independence TAR graph is unique.
Hamiltonicity and connectedness results for independent set reconfiguration also parallel those for vertex cover reconfiguration. Let $\underline{\alpha}$ denote the minimum cardinality of a maximal independent set. The greatest $k_{0}$ such that $\mathfrak{I}_{k}(G)$ is connected for all $k\leq k_{0}$ is denoted by $\alpha_{0}(G)$, and the greatest $k$ such that $\mathfrak{I}_{k}(G)$ is connected is denoted by $\underline{\alpha_{0}}(G)$. The next example shows that strict inequality is possible in the bound $\underline{\alpha}(G)-1\geq\alpha_{0}$ (cf. Proposition 8.4). It could be derived from Example 6.3 for vertex covering, but is also easy to see directly.
Example 8.19.
The bipartite graph $K_{p,q}$ for $2\leq p\leq q$ with partite sets $A$ and $B$ has exactly $A$ and $B$ as its maximal independent sets. Then $\alpha(K_{p,q})=q$, $\underline{\alpha}(K_{p,q})=p$, and the vertex of $\mathfrak{I}(K_{p,q})$ corresponding to the empty set is a cut-vertex. Thus $\alpha_{0}(K_{p,q})=0<p-1$; in fact, $\alpha_{0}(K_{p,q})=\alpha(K_{p,q})+\underline{\alpha}(K_{p,q})-n$. Furthermore, $K_{p,q}$ has a Hamilton path but not a Hamilton cycle.
Remark 8.20.
Let $G$ be a graph. Since the independence and vertex cover numbers are complementary parameters, $\mathfrak{I}(G)$ has a Hamiltonian path (cycle) if and only if $\mathfrak{C}(G)$ has a Hamiltonian path (cycle) by Remark 8.12. Examples illustrating equality or strict inequality in a connectedness bound for vertex covering illustrate the same for independence. See Section 6.1 for more examples.
Recall that an even-hole-free graph is one with no induced even cycles. Proposition 8.22, which shows many graphs $G$ satisfy $\alpha_{0}(G)=\underline{\alpha}(G)-1$, follows immediately from Proposition 6.8, but to obtain it that way would distort the historical record and deny the authors of the original work on independent sets their due. Instead, we present the original independence results that are used in all the reconfiguration results on even-hole-free graphs discussed in this paper.
Theorem 8.21 is based on Theorem 7 in [30], which is stated for token jumping (TJ) reconfiguration. The translation between token jumping and TAR reconfiguration for indepemdent sets is established in Theorem 1 of [30] (which is the basis for Proposition 2.44).
Theorem 8.21.
[30] If $T_{1}$ and $T_{2}$ are independent sets of size $k$ in a graph $G$ and $G[(T_{1}\setminus T_{2})\cup(T_{2}\setminus T_{1})]$ is even-hole-free, then there is a path between independent sets $T_{1}$ and $T_{2}$ in $\mathfrak{I}_{k-1}(G)$.
Proposition 8.22.
Let $G$ be an even-hole-free graph. Then $\alpha_{0}(G)=\underline{\alpha}(G)-1$.
Proof.
Let $\ell=\underline{\alpha}(G)$ and let $T_{1},T_{2}\in V(\mathfrak{I}_{\ell-1}(G))$. Then there exist maximal independent sets $M_{1}$ and $M_{2}$ such that $T_{i}\subseteq M_{i},i=1,2$. Without loss of generality, assume $|M_{1}|\geq|M_{2}|$ and let $k=|M_{2}|$; note that $k\geq\ell$. If $|T_{1}|>k$, choose $M^{\prime}_{1}\subseteq T_{1}$ such that $|M^{\prime}_{1}|=k$; otherwise, Choose $M^{\prime}_{1}$ such that $T_{1}\subseteq M^{\prime}_{1}$ and $|M^{\prime}_{1}|=k$. Then by Theorem 8.21, there exists a path between $M^{\prime}_{1}$ and $M_{2}$ in $\mathfrak{I}_{k-1}(G)$. As in the proof of Proposition 6.8, there is there is a path from $T_{1}$ to $T_{2}$ in $\mathfrak{I}_{\ell-1}(G)$. Thus $\mathfrak{I}_{\ell-1}(G)$ is connected, and $\alpha_{0}(G)=\underline{\alpha}(G)-1$ by Propositions 8.6(2) and 8.4(3). ∎
8.4 Irredundance and zero forcing irredundance
In this section we apply base graph complement parameter results to irredundance number and (standard) zero forcing irredundance number, which are defined in Section 1.2. The irredundance TAR graph of a base graph $G$ is denoted by $\mathfrak{Ir}(G)$ and the zero forcing irredundance TAR graph of a base graph $G$ is denoted by $\mathfrak{Zir}(G)$. The lower irredundance number of $G$ is denoted by $\operatorname{ir}(G)$ and the lower zero forcing irredundance number of $G$ is denoted by $\operatorname{zir}(G)$ (these follow the literature rather than our underline convention).
Proposition 8.23.
The upper irredundance number $\operatorname{IR}$ and the upper zero forcing irredundanace number $\operatorname{ZIR}$ are robust $Y$-set parameters.
Proof.
Let $G$ be a graph.
(I): Any subset of an Ir-set or ZIr-set of $G$ is an Ir-set or ZIr-set of $G$, respectively. Thus $\operatorname{IR}$ and $\operatorname{ZIR}$ are sub $Y$-set parameters.
(II): For $v\in V(G)$, $N[v]$ is a private neighborhood of $v$ relative to $\{v\}$, so $\{v\}$ is an Ir-set. Since $V(G)$ is a fort, it is a private fort of any one vertex $v$, and $\{v\}$ is a ZIr-set.
(III): Now assume $G_{1},\dots,G_{k}$ are the connected components of $G$. For any vertex $v\in V(G)$, there is some $i$ such that $v\in V(G_{i})$ and thus $N_{G}[v]\subseteq V(G_{i})$. Let $F$ be a fort. Then $F\cap V(G_{i})$ is a fort of $G_{i}$ if and only if $F\cap V(G_{i})$ is nonempty. Thus $T$ is a Ir-set of $G$ (respectively, a ZIr-set of $G$) if and only if $T\cap V(G_{i})$ is an Ir-set of $G$ (respectively, a ZIr-set of $G_{i}$) for $i=1,\ldots,k$. ∎
By Proposition 8.23, the main isomorphism results hold for $\mathfrak{Ir}(G)$ and $\mathfrak{Zir}(G)$ (note that $\operatorname{IR}(K_{1})=1$ and $\operatorname{ZIR}(K_{1})=1$).
Theorem 8.24.
Let $G$ and $G^{\prime}$ be graphs. If $\mathfrak{Ir}(G)\cong\mathfrak{Ir}(G^{\prime})$, then $G$ and $G^{\prime}$ have the same order and there is a relabeling of the vertices of $G^{\prime}$ such that $G$ and $G^{\prime}$ have exactly the same Ir-sets.
Theorem 8.25.
Let $G$ and $G^{\prime}$ be graphs. If $\mathfrak{Zir}(G)\cong\mathfrak{Zir}(G^{\prime})$, then $G$ and $G^{\prime}$ have the same order and there is a relabeling of the vertices of $G^{\prime}$ such that $G$ and $G^{\prime}$ have exactly the same ZIr-sets.
Note that $\operatorname{IR}$ and $\operatorname{ZIR}$ are not the base graph complement parameters to domination and standard zero forcing, even though we have used base graph complementation to establish the previous two results. Next we discuss some examples of irredundance TAR graphs and zero forcing irredundance TAR graphs, including graphs with unique TAR graphs and nonunique TAR graph.
Remark 8.26.
Let $G$ be a graph on $n$ vertices. It is immediate that $\operatorname{ir}(G)=\operatorname{IR}(G)=n$ if and only if $G=\overline{K_{n}}$. By [12, Remark 5.1], $\operatorname{zir}(G)=\operatorname{ZIR}(G)=n$ if and only if $G=\overline{K_{n}}$. Thus, $\mathfrak{Ir}(\overline{K_{n}})$ and $\mathfrak{Zir}(\overline{K_{n}})$ are unique by Theorems 8.24 and 8.25. Moreover, $\mathfrak{Ir}(\overline{K_{n}})\cong\mathfrak{Zir}(\overline{K_{n})}\cong Q_{n}$ and therefore $\mathfrak{Ir}(\overline{K_{n}})$ and $\mathfrak{Zir}(\overline{K_{n}})$ have Hamiltonian cycles.
Similarly, let $G$ be a graph of order $n$ with no isolated vertices. If $u$ and $v$ are not adjacent, then $\{u,v\}$ is an IR-set, so $\operatorname{IR}(G)=1$ if and only if $G\cong K_{n}$. Thus, $\mathfrak{Ir}({K_{n}})$ is unique by Theorem 8.24. Moreover, $\mathfrak{Ir}({K_{n})}\cong K_{1,n}$ and hence does not have a Hamiltonian path for $n\geq 3$. By [12, Remark 5.2], $\operatorname{zir}(G)=\operatorname{ZIR}(G)=n-1$ if and only if $G={K_{n}}$. Thus, $\mathfrak{Zir}({K_{n}})$ is unique by Theorem 8.25. Moreover, $\mathfrak{Zir}({K_{n})}\cong Q_{n}-v$ and therefore has a Hamiltonian path but not a Hamiltonian cycle.
Proposition 8.27.
Let $G$ be a graph and $n\geq 1$. If $\mathfrak{Zir}(G)\cong\mathfrak{Zir}(P_{n})$, then $G\cong P_{n}$. If $\mathfrak{Zir}(G)\cong\mathfrak{Zir}(K_{1,n-1})$, then $G\cong K_{1,n-1}$.
Proof.
Assume $\mathfrak{Zir}(G)\cong\mathfrak{Zir}(P_{n})$ or $\mathfrak{Zir}(G)\cong\mathfrak{Zir}(K_{1,n-1})$. In [12] it is shown that $\operatorname{zir}(G)=1$ if and only if $G\cong P_{n}$ or $G\cong K_{1,n-1}$. By Theorem 8.24, $G\cong P_{n}$ or $G\cong K_{1,n-1}$. Note that if $n\leq 3$, then $P_{n}\cong K_{1,n-1}$, so it suffices to show that $\mathfrak{Zir}(P_{n})\not\cong\mathfrak{Zir}(K_{1,n-1})$ for $n\geq 4$.
Let $v_{1}$ and $v_{n}$ be the vertices of degree 1 in $P_{n}$. Then every fort of $P_{n}$ contains both $v_{1}$ and $v_{n}$. Thus, $\{v_{1}\}$ and $\{v_{n}\}$ are maximal ZIr-sets and hence have degree 1 in $\mathfrak{Zir}(P_{n})$.
Let $v$ be the vertex of degree $n-1$ in $K_{1,n-1}$. The only fort that contains $v$ is $V(K_{1,n-1})$. Thus, $\{v\}$ is a maximal ZIr-set and has degree 1 in $\mathfrak{Zir}(K_{1,n-1})$. Note that every pair of vertices in $V(G)\setminus\{v\}$ is a fort of $K_{1,n-1}$. Thus for $n\geq 4$, $\{v\}$ is the only maximal ZIr-set of size 1. It follows that $\mathfrak{Zir}(P_{n})\not\cong\mathfrak{Zir}(K_{1,n-1})$ for $n\geq 4$. ∎
The next example is motivated by [7, Example 3.3].
Example 8.28.
Let $n\geq 4$. Then $\mathfrak{Zir}(C_{n})=\mathfrak{Zir}(C_{n}+vw)$ for any two vertices $v$ and $w$ that are not adjacent in $C_{n}$. A set $S$ is standard zero forcing set if and only it it contains two vertices that are adjacent on the cycle; a set that does not contain two such vertices cannot perform any force. Thus, $C_{n}$ and $C_{n}+vw$ have the same forts and so $\mathfrak{Zir}(C_{n})=\mathfrak{Zir}(C_{n}+vw)$.
8.5 Hypercube complement parameters
In this section we introduce a different way to convert an $X$-set parameter into a $Y$-set parameter by taking the complement of the vertex set of $\mathfrak{X}(G)$ in the set of hypercube vertices (the power set of the set of vertices of the base graph $G$). These results are then applied to failed zero forcing in the next section.
Definition 8.29.
(Hypercube complement parameters)
Let $X$ be a super $X$-set parameter such that $\emptyset$ is not an $X$-set of $G$ for any graph $G$. We say that a subset $T\subseteq V(G)$ is an $Y_{H(X)}$-set if and only if $T$ is not a $X$-set of $G$, and we define the parameter $Y_{H(X)}(G)$ to be the maximum cardinality of a $Y_{H(X)}$-set of $G$.
Let $Y$ be a sub $Y$-set parameter such that $V(G)$ is not an $Y$-set of $G$ for any graph $G$. We say that a subset $S\subseteq V(G)$ is an $X_{H(Y)}$-set if and only if $S$ is not a $Y$-set of $G$, and we define the parameter $X_{H(Y)}(G)$ to be the minimum cardinality of an $X_{H(Y)}$-set of $G$.
The parameters $X$ and $Y_{H(X)}$ (or $Y$ and $X_{H(Y)}$) are called hypercube complement parameters. The $X_{H(Y)}$-TAR and $Y_{H(X)}$-TAR graphs of a base graph $G$ are denoted by ${\mathfrak{X}}_{H(Y)}(G)$ and ${\mathfrak{Y}}_{H(X)}(G)$, respectively. The lower $Y_{H(X)}$ number of $G$ is denoted by $\underline{Y}_{H(X)}(G)$ and the upper $X_{H(Y)}$ number of $G$ is denoted by $\overline{X}_{H(Y)}(G)$
Remark 8.30.
The following statements are immediate from the definitions (because the assumptions $\emptyset$ is not an $X$-set and $V(G)$ is not an $Y$-set guarantee there is an $Y_{H(X)}$-set and $X_{H(Y)}$-set for every graph. If $X$ is a super $X$-set parameter, then $Y_{H(X)}$ is a sub $Y$-set parameter. If $Y$ is a sub $Y$-set parameter, then $X_{H(Y)}$ is a super $X$-set parameter. The $Y_{X_{H(Y)}}$-sets of $G$ are the $Y$-sets of $G$ and $Y_{X_{H(Y)}}(G)=Y(G)$ and analogous statements hold when $Y$ and $X$ are interchanged.
The definition of hypercube graph complement parameters is illustrated in the next example.
Example 8.31.
Number the vertices of the path $P_{4}$ in path order. The minimal standard zero forcing sets of $P_{4}$ are $\{1\},\{4\}$ and $\{2,3\}$, and $\mathfrak{Z}(P_{4})$ is shown in Figure 8.2(a) with the part of the hypercube of all subsets that are not $\operatorname{Z}$-sets shown in gray. The maximal $Y_{H(\operatorname{Z})}$-sets of $P_{4}$ are $\{2\}$ and $\{3\}$, and ${\mathfrak{Y}}_{H(\operatorname{Z})}(P_{4})$ is shown in Figure 8.2(b).
$\{1,2,3,4\}$$\{1,2,3\}$$\{1,2,4\}$$\{1,3,4\}$$\{2,3,4\}$$\{1,2\}$$\{1,3\}$$\{2,3\}$$\{1,4\}$$\{2,4\}$$\{3,4\}$$\{1\}$$\{2\}$$\{3\}$$\{4\}$$\emptyset$(a)$\{1,2,3,4\}$$\{1,2,3\}$$\{1,2,4\}$$\{1,3,4\}$$\{2,3,4\}$$\{1,2\}$$\{1,3\}$$\{1,4\}$$\{2,3\}$$\{2,4\}$$\{3,4\}$$\{1\}$$\{2\}$$\{3\}$$\{4\}$$\emptyset$(b)
Figure 8.2: The graphs $\mathfrak{Z}(P_{4})$ and ${\mathfrak{Y}}_{H(\operatorname{Z})}(P_{4})$, illustrating the isomorphism
It is not true that the hypercube complement of a robust $X$-set parameter must be robust (see Example 8.34). However, if $X$ is a robust $X$-set parameter, then $Y_{H(X)}$ satisfies Theorem 2.23.
Theorem 8.32.
Let $X$ be a robust $X$-set parameter and let $G$ and $G^{\prime}$ be base graphs such ${\mathfrak{Y}}_{H(X)}(G)\cong{\mathfrak{Y}}_{H(X)}(G^{\prime})$. If $X(K_{1})=0$ or $G$ and $G^{\prime}$ have no isolated vertices, then $G$ and $G^{\prime}$ have the same order and there is a relabeling of the vertices of $G^{\prime}$ such that $G$ and $G^{\prime}$ have exactly the same $Y_{H(X)}$-sets.
Proof.
It follows from the definition of the hyercube complement parameter that $\mathfrak{X}(G)\cong\mathfrak{X}(G^{\prime})$. Assume $X(K_{1})=0$ or $G$ and $G^{\prime}$ have no isolated vertices. Then by Theorem 2.23, $G$ and $G^{\prime}$ have the same order and there is a relabeling of the vertices of $G^{\prime}$ such that $G$ and $G^{\prime}$ have exactly the same $X$-sets. This implies $G$ and $G^{\prime}$ have exactly the same $Y_{H(X)}$-sets by the definition of the hyerpcube complement parameter ∎
All the elementary connectedness results for sub $Y$-set parameters apply to the hypercube complement parameter $Y_{H(X)}$ of a super $X$-set parameter, including Propositions 8.4 and 8.6 and Corollary 8.5.
8.6 Failed zero forcing
In this section we apply hypercube complement parameter results to the failed standard zero forcing number, which is defined in Section 1.2. The failed zero forcing number TAR reconfiguration graph of a base graph $G$ is denoted by ${\mathfrak{F}}(G)$ and the lower failed zero forcing number of $G$ is denoted by $\underline{F}(G)$.
Remark 8.33.
Failed zero forcing is the hypercube complement of zero forcing. That is, $\operatorname{F}(G)=Y_{H(\operatorname{Z})}(G)$, $\underline{F}(G)=\underline{Y}_{H(\operatorname{Z})}(G)$, and ${\mathfrak{F}}(G)={\mathfrak{Y}}_{H(\operatorname{Z})}(G)$.
The next example shows failed zero forcing is not a robust $Y$-set parameter.
Example 8.34.
Note that for a path $P_{n}$ with end vertices $v_{1}$ and $v_{n}$, any set that contains $v_{1}$ or $v_{n}$ is a zero forcing set and thus the singletons $\{v_{1}\}$ and $\{v_{n}\}$ are not failed zero forcing sets. Therefore, $F$ is not a robust $Y$-set parameter. Since failed zero forcing is the hypercube complement of zero forcing, Figure 8.2(b) shows ${\mathfrak{F}}(P_{4})$.
The next result now follows from the analogous result for hypercube complement parameters (and the fact that the zero forcing number is a robust $X$-set parameter)
Theorem 8.35.
Let $G$ and $G^{\prime}$ be base graphs such ${\mathfrak{F}}(G)\cong{\mathfrak{F}}(G^{\prime})$. Then $G$ and $G^{\prime}$ have the same order and there is a relabeling of the vertices of $G^{\prime}$ such that $G$ and $G^{\prime}$ have exactly the same failed zero forcing sets.
9 Concluding remarks
In Section 2 we have established that with a few exceptions, the results for original $X$-set parameters in [6] and [7] remain valid for robust $X$-set parameters (with the same proofs). Connectedness results are central to the study of reconfiguration (transforming one solution to another with every intermediate intermediate step being a solution). Although connectedness examples tend to be very parameter specific, in Section 2.5 most connectedness results established in [6] for original $X$-set parameters are extended to super $X$-set parameters.
Results for specific parameters are surveyed and extended in Sections 3-6. For many of these parameters, there are open questions remaining to investigate. For example, we do not have an example of a PSD TAR graph that has a Hamilton cycle.
We introduce the universal perspective on TAR reconfiguration for maximizing parameters (sub $Y$-set parameters) in Section 8, and many avenues remain unexplored. What does simplical complex theory tell us about robust $Y$-set parameters? We introduced two ways of complementing a sub $Y$-set parameter to obtain a super $X$-set parameter and vice versa, base graph complements and hypercube complements. There may be additional useful ways to connect sub $Y$-set parameters and super $X$-set parameters. There is more work to be done on TAR reconfiguration of specific sub $Y$-set parameters.
Complexity of reconfiguration problems has been studied extensively (see, for example, [28, 30, 32, 34]) but is not surveyed here. Most of this work focuses on the complexity of reconfiguration of specific parameters. It would be interesting to develop a universal theory of reconfiguration complexity for $X$-set and $Y$-set parameters; we hope the work here will be useful for this.
Acknowledgements
The research of all authors was partially supported by NSF grant 2331634. The research of B. Curtis is also partially supported by NSF grant 1839918.